### v1

# Improving exponential-family random graph models for bipartite networks

Abstract
Bipartite graphs, representing two-mode networks, arise in many research fields. These networks have two disjoint node sets representing distinct entity types, for example persons and groups, with edges representing associations between the two entity types. In bipartite graphs, the smallest possible cycle is a cycle of length four, and hence four-cycles are the smallest structure to model closure in such networks. Exponential-family random graph models (ERGMs) are a widely used model for social, and other, networks, including specifically bipartite networks. Existing ERGM terms to model four-cycles in bipartite networks, however, are relatively rarely used. In this work we demonstrate some problems with these existing terms to model four-cycles, and define new ERGM terms to help overcome these problems. The position of the new terms in the ERGM dependence hierarchy, and their interpretation, is discussed. The new terms are demonstrated in simulation experiments, and their application illustrated with ERGM models of empirical networks ranging in size from hundreds of nodes to hundreds of thousands of nodes.
Keywords— bipartite graph, two-mode network, exponential-family random graph model, ERGM, four-cycle
1 Introduction
Bipartite graphs are graphs whose nodes can be partitioned into two disjoint sets, such that an edge exists only between nodes in different sets. Such graphs have important applications in representing two-mode networks, which are networks in which there are two types of nodes, with edges possible only between nodes of different types. An important example of a two-mode network is an affiliation network, in which one type of node represents a person, the other type of node represents a group, and an edge represents membership of a person in a group [1]. Two-mode networks have applications not only in sociology, but also biology, ecology, political science, psychology, finance, and economics; for a recent review of applications and methods for two-mode networks, see Neal et al. [2]. Bipartite networks also arise as representing the meso-level network in the conceptualization and analysis of multilevel networks [3].
Two-mode networks can be studied by means of their projections onto one-mode (unipartite) networks, thereby allowing the use of existing methods for one-mode networks, however this can result in lost information, and properties of the one-mode networks (such as high clustering coefficients) that are due to the projection process rather than the original data [4]. Although the former problem can be overcome by using both projections in analyses [5, 6], it is still desirable to directly study the original two-mode network, for which specific methods are required [4].
In studying one-mode networks, a central concept is triadic closure, the tendency for a path of length two (a “two-path”; three nodes connected by two edges) to be “closed” into a triangle by the addition of a third edge. In the context of a social network, this is the process of a friend of a friend becoming themselves a friend, and is perhaps most well known via the “strength of weak ties” [7] argument, whereby an open two-path of strong ties is the “forbidden triad”, which is “forbidden” because the two actors with strong ties to a common third actor must themselves have a strong tie. In a bipartite network, however, a closed triad (triangle) is impossible; indeed it is a defining feature of bipartite graphs that only cycles of even length are possible [8]. Therefore, the smallest possible cycle in a bipartite graph is a four-cycle, and hence four-cycles are frequently used to measure closure in bipartite networks [9, 10].
Exponential random graph models (ERGMs) are a widely used model for social [11, 12, 13, 14], and other [15, 16, 17] networks. Specific forms of the ERGM have been developed for two-mode networks [18, 19, 20, 21], however, as we shall show in this work, existing ERGM models for bipartite networks often have problems modelling four-cycles, and hence can frequently not adequately model closure in bipartite networks.
In this work, we will show that, despite the importance of four-cycles in two-mode networks, ERGM terms to model four-cycles in such networks are relatively rarely used, in contrast to the ubiquity of terms modelling triadic closure in one-mode networks. We will then describe some problems with existing configurations for modelling four-cycles that could explain their relatively infrequent use, and propose new ERGM configurations for modelling four-cycles to help overcome these problems. We will discuss the interpretation of these new parameters, and their position in the dependence hierarchy of Pattison & Snijders [22]. We will then demonstrate the new configurations using simulation experiments and apply them in ERGM modelling of seven empirical networks in a variety of domains (ecology, criminology, and sociology) and ranging in size from hundreds of nodes to hundreds of thousands of nodes.
2 The exponential-family random graph model (ERGM)
An ERGM is a probability distribution with the form
$\Pr_{\theta}(X=x)=\frac{1}{\kappa(\theta)}\exp\left(\sum_{C}\theta_{C}g_{C}(x)\right)$ (1)
where
  • $X=[X_{ij}]$ is square binary matrix of random tie variables,
  • $x$ is a realization of $X$,
  • $C$ is a “configuration”, a set of nodes and a subset of ties between them, designed in order to model a particular structure of interest,
  • $g_{C}(x)$ is the network statistic for configuration $C$,
  • $\theta$ is a vector of model parameters, where each $\theta_{C}$ is the parameter corresponding to configuration $C$,
  • $\kappa(\theta)=\sum_{x\in G_{n}}\exp\left(\sum_{C}\theta_{C}g_{C}(x)\right)$, where $G_{n}$ is the set of all square binary matrices of order $n$ (graphs with $n$ nodes), is the normalising constant to ensure a proper distribution.
We will use the notation $x_{ij}$ (where $1\leq i\leq n$, $1\leq j\leq n$) for elements of the binary adjacency matrix $x$. In this work we consider only the case of undirected networks, so $x_{ij}=x_{ji}$, and the cardinality of $G_{n}$ is $\lvert G_{n}\rvert=2^{n(n-1)/2}$. In the case of bipartite networks, the two disjoint node sets are denoted $A$ and $B$, with sizes $N_{A}=\lvert A\rvert$ and $N_{B}=\lvert B\rvert$ respectively, and so $N_{A}+N_{B}=n$, and $x_{ij}=0$ if both $i$ and $j$ are in node set $A$ or both $i$ and $j$ are in node set $B$. In this bipartite case, the normalising constant is $\kappa(\theta)=\sum_{x\in G^{\mathrm{Bipartite}}_{N_{A},N_{B}}}\exp\left(\sum_{C}\theta_{C}g_{C}(x)\right)$, where $G^{\mathrm{Bipartite}}_{N_{A},N_{B}}$ is the set of all bipartite graphs with node set sizes $N_{A}$ and $N_{B}$. This set has cardinality $\lvert G^{\mathrm{Bipartite}}_{N_{A},N_{B}}\rvert=2^{N_{A}N_{B}}$.
Estimating the value of the parameter vector $\theta$ which maximises the probability of the observed graph, that is, the maximum likelihood estimator (MLE), enables inferences regarding the under-representation (negative and statistically significant estimate) or over-representation (positive and statistically significant estimate) of the corresponding configurations. These inferences are conditional on the other configurations included in the model, which need not be independent.
Estimating the MLE of (1) is computationally intractable due to the normalising constant $\kappa(\theta)$ (specifically, the size of the set of graphs it sums over). Therefore, Markov chain Monte Carlo (MCMC) methods are usually used [23, 24, 25]. One such algorithm is the “Equilibrium Expectation” (EE) algorithm [26, 27, 28], which was recently shown to converge to the MLE [28, 17].
ERGM models with only simple configurations (such as the Markov edge plus triangle model) can be prone to problems with phase transitions or “near-degeneracy” [29, 30, 31, 32, 33, 34, 35, 14]. Such problems are typically avoided by the use of “alternating” [30, 36, 37] or “geometrically weighted” [38, 39] configurations. The former are parameterised with a decay parameter $\lambda$ controlling the rate at which the weight of contributions from additional terms in the statistic decay. The corresponding parameter for the geometrically weighed configurations can be estimated as part of the model, in which case it becomes a “curved ERGM” [40], however in this work we will use fixed values of $\lambda$ for the “alternating” configurations.
3 Literature survey
In order to get an overview of which effects are used in modelling bipartite networks with ERGMs, and how well they fit four-cycles, we conducted a comprehensive survey of publications which included ERGM models of bipartite networks. To be included, a publication must contain one or more ERGM models of one or more empirical bipartite (two-mode) networks. There must be sufficient detail given to know at least the parameters included in the models, and their estimated signs and statistical significance. Models of one-mode projections of two-mode networks were excluded; we only consider ERGM models of the bipartite network itself. Models of multilevel networks were excluded, although if there is a model of just the cross-level (bipartite) network, this is included.
We used Google Scholar (search date 24 November 2022) to search for papers with terms “bipartite“ and “ergm”,“bpnet” and “ergm”, “statnet” and “bipartite”, and “gwb1dsp” or “gwb2dsp” (the latter being statnet [41, 42] terms used for bipartite networks). We also included papers that we had prior knowledge of being about bipartite ERGMs, and followed backward and forward citations (using Google Scholar for the latter) to find other relevant papers. The papers included in the survey are listed in Table LABEL:tab:literature (Appendix A).
Most of the models in Table LABEL:tab:literature were estimated with the BPNet [43], MPNet [44, 45] or statnet [41, 46, 47, 42, 48, 49, 50] software, but a small number were estimated either by maximum pseudo-likelihood estimation (MPLE) or with Bayesian methods using the Bergm [51, 52] software. The one model included that was estimated with Bergm contains no terms to model four-cycles or goodness-of-fit tests including four-cycles [53]. Of the ten models (across three publications) estimated by MPLE, only one contains a term to model four-cycles, and this is found to be positive and significant [54].
Table 1 summarises the parameter estimates in models from the literature in Table LABEL:tab:literature that were estimated using BPNet or MPNet, and contain the four-cycles parameter C4, or the bipartite alternating $k$-two-path parameters ($\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ in BPNet, XACA and XACB in MPNET) defined in Wang et al. [18]. Less than a third ($20/63$) include the four-cycles parameter, less than half ($30/63$) include either of the two alternating $k$-two-path parameters, and less than 30% ($17/63$) include both $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$.
Table 1: Counts of parameters in models estimated by BPNet or MPNet (total 63) in the reviewed literature.
C4 $\mathrm{K\text{-}C_{P}}$ $\mathrm{K\text{-}C_{A}}$
Total estimated 20 30 25
Negative 14 20 15
Negative and signif. 5 14 9
Positive 6 10 10
Positive and signif. 4 7 2
Table 2 summarises the parameter estimates in models from the literature in Table LABEL:tab:literature that were estimated using statnet, and which contain the four-cycle term, the bipartite geometrically weighted dyadwise shared partner distribution (gwb1dsp or gwb2dsp) term, the statnet equivalent of the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ parameters, or the geometrically weighted non-edgewise shared partner (gwnsp) term. The counts for gwb1dsp and gwb2dsp include those for the bipartite geometrically weighted non-edgewise shared partner terms gwb1nsp and gwb2nsp, which are equivalent to gwb1dsp and gwb2dsp, respectively, for bipartite networks. Only one model estimated with statnet contained an explicit term for four-cycles [55]. Further, only $5/43$ models contain a gwb1dsp pr gwb2dsp term at all, and only two models contain both (Lubell & Robbins [56] have two models, both of which include both gwb1nsp and gwb2nsp).
Table 2: Counts of parameters in models estimated by statnet (total 43) in the reviewed literature. The counts for gwb1dsp and gwb2dsp include those for the equivalent parameters gwb1nsp and gwb2nsp, respectively.
cycle(4) gwb1dsp gwb2dsp gwnsp
Total estimated 1 2 5 4
Negative 1 1 2 1
Negative and signif. 1 1 0 0
Positive 0 1 3 3
Positive and signif. 0 1 2 3
Less than a quarter ($26/117$) of the models include an explicit assessment of goodness-of-fit to four-cycles, and of those, the majority ($18/26$) are good. Only seven of these are for models that explicitly include C4 as a model parameter, and, as expected of any converged model containing this term, these models fit four-cycles well. Conversely, all of the models which are described as having a poor fit to four-cycles in the goodness-of-fit procedure are models that do not contain the C4 parameter. However, of these eight models, five contain either $\mathrm{K\text{-}C_{A}}$ or $\mathrm{K\text{-}C_{P}}$, and one contains both.
This relative rarity of models containing terms to model closure (four-cycles) in bipartite networks, or assess goodness-of-fit to four-cycles, is in stark contrast to ERGM modelling for one-mode networks, where terms modelling triadic closure, such as triangles, alternating-$k$-triangles, or geometrically weighted edgewise shared partners (gwesp in statnet) are almost always included in models, since triadic closure (as evidenced by clustering, or transitivity, in the network), is a well-known feature of social networks [7, 57, 58]. For example, Clark & Handcock [59] use the latent order logistic (LOLOG) model [60] to reproduce ERGM models for 13 networks from peer reviewed papers, and all of these models contain alternating-$k$-triangles [36, 61, 62, 63], gwesp [64, 65, 66, 67, 68, 69], or three-cycles [61, 70] terms, and in the majority of cases the results are able to be replicated with a triangle term in the LOLOG model [59]. And yet, despite higher than expected numbers of four-cycles (or bipartite clustering coefficient) being a notable feature of some bipartite networks, such as director interlock networks [9, 10], and collaboration and communication networks [10], the literature survey presented here shows that terms to model four-cycles are relatively rarely used in published ERGM models of bipartite networks. This could be due to researchers choosing not to model bipartite closure, however, as shown by some examples in the following section, it can also be due to difficulties in obtaining converged model estimates when using existing model terms.
4 Problems with existing bipartite ERGM statistics
It is notable that in the original paper proposing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ configurations [18], they are explicitly described as $k$-two-path statistics, and in fact they are just bipartite versions of the one-mode alternating-$k$-two-path statistic, representing multiple shared partners (gwdsp in statnet, and then later gwb1dsp and g2b2dsp for bipartite networks). There is, however, already some “semantic slippage” into interpreting them as “cycles” or “closure” — even though they only actually include cycles (closure) when $k>1$ ($k=1$ is a two-path, $k=2$ is a four-cycle; see Fig. 3 in Wang et al. [18]). For example, “… a better chance of achieving model convergence when closure effects ($K-C_{P}$ and $K-C_{A}$) are included in the model.” [18, p. 22]. Even the name $\mathrm{K\text{-}C_{A}}$ (or $\mathrm{K\text{-}C_{P}}$) suggests “cycle” or “closure” by the use of the “C” (C4 is used for four-cycles in the paper). Wang et al. [19] goes back to naming the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ statistics as A2P-A and A2P-B and describing them as “shared affiliations (alternating two-paths)” [19, p. 213]. However the MPNet terminology is XACA and XACB [44, 45], again with the “C” suggestive of cycles or closure, and the EstimNetDirected [71] software (https://github.com/stivalaa/EstimNetDirected) refers to these effects as BipartiteAltKCyclesA and BipartiteAltKCyclesB.
In the influential book edited by Lusher et al. [11], Wang [20] describes $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ explicitly as “alternating A cycles” and “alternating P cycles” [20, p. 124], with [20, Fig. 10.11] captioned “Alternating 2-paths” but with the figure panels labelled “A cycles (KCA)” and “P cycles (KCP)” respectively [20, p. 125]. This description or interpretation is carried over into the empirical part of the book, with Harrigan & Bond [72], applying bipartite ERGM to a director interlock network, describing $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ as “alternating $k$-cycles” for directors and corporations respectively [72, p. 266], and in the ERGM results tables as “Director 4-cycles” and “Corporation 4-cycles” [72, pp. 268-269]. It is also notable that Harrigan & Bond [72] describes the difficulty of fitting models with $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ and the resulting poor goodness-of-fit for the four-cycle statistic:
We found for this particular network that we cannot have both K-Cp and K-Ca in the same model due to convergence issues. In line with Wang, Sharpe, Robins, and Pattison (2009), we present two alternative models, one with each possible k-cycle parameter. … However, in line with Wang, Sharpe, Robins, and Pattison (2009), we had a poor fit on the classic 4-cycle parameter [C4], which suggests that improving these structural effects is a substantial area of future research.
[72, p. 267]
Given the results of the literature review described in Section 3 (which includes Harrigan & Bond [72]) it would seem, however, that no such research improving these structural effects has been published yet, and this work may be the first attempt to do so.
We may conclude from this that the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics count too many things other than four-cycles to be usefully used and interpreted as bipartite closure in many cases. Specifically, they count two-paths as their first (highest weighted) term. This results in situations where long paths or cycles contribute to the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ statistics, despite having exactly zero four-cycles (see Table 3). Perhaps even more problematically, high-degree nodes, or stars, for example the “Nine-star” structure in Table 3, result in large values of $\mathrm{K\text{-}C_{A}}$ or $\mathrm{K\text{-}C_{P}}$ (XACA or XACB in MPNet terminology), depending on which node set the hub node is in, but also have exactly zero four-cycles. It seems clear that large networks are likely to contain many stars, and many paths (of length two or more; and note that a two-star is just a two-path), as well as large cycles [73, 74], and these will contribute to large values of the $\mathrm{K\text{-}C_{P}}$ and/or $\mathrm{K\text{-}C_{A}}$ statistics, but nothing to the number of four-cycles (the C4 statistic).
Table 3: Statistics of some example bipartite networks. L is the number of edges, C4 the number of four-cycles, and BpNP4CA and BpNP4CB are the new statistics BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB. Nodes in node set A are represented as red circles, and nodes in node set B as blue squares.
Name Visualization $\mathrm{N_{A}}$ $\mathrm{N_{B}}$ L C4 XACA XACB BpNP4CA BpNP4CB
Two-path [Uncaptioned image] 1 2 2 0 1 0 0 0
Four-cycle [Uncaptioned image] 2 2 4 1 1.5 1.5 2 2
Four-cycles-3 [Uncaptioned image] 2 3 6 3 4.5 1.75 3.4641 4.24264
Ten-cycle [Uncaptioned image] 5 5 10 0 5 5 0 0
Nine-star [Uncaptioned image] 1 9 9 0 36 0 0 0
Four-fan-3 [Uncaptioned image] 4 6 12 3 16.5 4.5 4.73205 6
In addition, interpretation of the alternating $k$-two-paths parameter is problematic: “In this article, we do not concentrate on alternating $k$-two-path parameters. For some data, we have found it important to include them in models but further work is needed to understand better their effect when included with other parameters” [36, p. 201]. Martin [75] described this as the authors “being somewhat mystified by this statistic” [75, p. 86], and that most ERGM modellers would not be able to describe a “clear behavioral-process analogue to the once-canonical alternating two paths statistic” [75, p. 86].
Therefore, we propose a new statistic that counts four-cycles (and not two-paths), but which is less prone to problems with near-degeneracy than the simple four-cycles parameter. (A simpler new statistic, based on $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ but which does not count two-paths, which turns out to be more prone to near-degeneracy than $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$, is described in Appendix B).
5 New statistics for modelling four-cycles in bipartite ERGMs
Let
$L_{2}(i,j)=\sum_{h\neq i,j}x_{ih}x_{hj}$ (2)
be the number of two-paths connecting node $i$ to node $j$. Since a four-cycle is a combination of two two-paths [30, p. 123], the number of four-cycles is
$C_{4}=\frac{1}{2}\sum_{i<j}\binom{L_{2}(i,j)}{2}.$ (3)
The sum in equation (3) is over the $\binom{n}{2}$ pairs of nodes in the graph, with the factor of $\frac{1}{2}$ to account for the double-counting due to the symmetry of each four-cycle containing two distinct pairs of nodes, each connected by two two-paths. The number of four-cycles containing a particular node $i$ is
$C_{4}(i)$ $=\sum_{j\neq i}\binom{L_{2}(i,j)}{2}$ (4)
=$=\sum_{\left\{j\,:\,d(i,j)=2\right\}}\binom{L_{2}(i,j)}{2}$ (5)
where $d(i,j)$ is the geodesic distance from $i$ to $j$. Some illustrative examples of the value of $C_{4}(i)$ for different nodes in some small graphs are shown in Fig. 1.
Figure 1: $C_{4}(i)$ is the number of four-cycles involving a node $i$.
The total number of four-cycles (3) can also be expressed in terms of the number of four-cycles at each node (4) as
$C_{4}=\frac{1}{4}\sum_{i}C_{4}(i)$ (6)
where the factor of $\frac{1}{4}$ accounts for the fact that each four-cycle is counted four times, once for each node it contains. The FourCyclesNodePower statistic is then defined as
$z_{\text{FourCyclesNodePower}(\alpha)}=\sum_{i}\left[C_{4}(i)^{\alpha}\right]$ (7)
where $0<\alpha\leq 1$ is the exponent for, in the terminology of Wilson et al. [76], the “$\alpha$-inside” weighting, since the subgraph counts ($C_{4}(i)$ here) are exponentiated before summing over all subgraphs. The “$\alpha$-outside” weighting would be to exponentiate the statistic after summing over all subgraphs, that is, in this case it would be $\left[\sum_{i}C_{4}(i)\right]^{\alpha}$. As discussed in Wilson et al. [76], the $\alpha$-inside weighting leads to local dependence as usually used in ERGMs, while the $\alpha$-outside weighting leads to global dependence, in which all ties are dependent on each other to some degree [76, p. 41]. The nature of the local dependencies induced by the new change statistics defined here using the $\alpha$-inside weighting are discussed in Section 5.2 below.
The statistics described in this section so far are equally applicable to one-mode and two-mode (bipartite) graphs. When dealing with bipartite graphs, however, it is often useful to consider statistics of the two node sets separately. Hence we also define
$z_{\text{BipartiteFourCyclesNodePowerA}(\alpha)}=\sum_{i\in A}\left[C_{4}(i)^{\alpha}\right]$ (8)
and
$z_{\text{BipartiteFourCyclesNodePowerB}(\alpha)}=\sum_{j\in B}\left[C_{4}(j)^{\alpha}\right]$ (9)
for the two node sets $A$ and $B$ respectively. Because the sets $A$ and $B$ are disjoint, we have
$z_{\text{FourCyclesNodePower}(\alpha)}=z_{\text{BipartiteFourCyclesNodePowerA}(\alpha)}+z_{\text{BipartiteFourCyclesNodePowerB}(\alpha)}.$ (10)
Representations of the new configurations are shown in Fig. 2. The vertical ellipsis $\vdots$ in the figures is to indicate that the configuration includes any number (up to $\lfloor(n-1)/3\rfloor$, since, apart from the one shared node, each four-cycle must include at most three distinct nodes) of four-cycles all involving a shared node (the central node in the figures).
Figure 2: Representations of the new configurations (left to right) FourCyclesNodePower, BipartiteFourCyclesNodePowerA, and BipartiteFourCyclesNodePowerB. Nodes in node set A are shown as red circles, and nodes in node set B are shown as blue squares.
Table 3 shows the values of the BipartiteFourCyclesNodePowerA (BpNP4CA) and BipartiteFourCyclesNodePowerB (BpNP4CB) statistics for some small example bipartite networks, with the parameter $\alpha=0.5$ (the alternating $k$-two-path statistics XACA and XACB are shown with their parameter $\lambda=2.0$). Note that, unlike the XACA and XACB ($\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$) statistics, these new statistics have the value zero for structures that contain no four-cycles (such as the two-path, ten-cycle, and nine-star structures in Table 3).
The “four-fan-3” graph in Table 3 is the same graph as the representation of the BipartiteFourCyclesNodePowerA configuration in Fig. 2 (ignoring the vertical ellipsis so that exactly three four-cycles are present). Note, however, that this has the counterintuitive property that, although it is a representation of the BipartiteFourCyclesNodePowerA configuration, in fact the value of the BipartiteFourCyclesNodePowerB statistic is greater than that of the BipartiteFourCyclesNodePowerA statistic for this graph. This is because the value of the statistic [(8) or (9)] is the sum over all nodes in the relevant node set ($A$ or $B$, respectively) of the four-cycle count at each node (4) raised to the power $\alpha$ (“$\alpha$-inside” weighting). Therefore, in this graph, the nodes in mode $B$ contribute more to the total as each one (of the six) is involved in exactly one four-cycle (and hence raising to the power of $\alpha$ still contributes one to the sum), while of the four nodes in mode $A$, three are involved in only one four-cycle, while the fourth is involved in three four-cycles and hence contributes only $3^{\alpha}\approx 1.73205$ (when $\alpha=0.5$). Generalising this four-fan-3 graph to four-fan-$k$ ($k\geq 1$, and if $k=1$ the graph is just a four-cycle) with the central high-degree node in node set $A$, we have:
$N_{A}$ $=k+1$ (11)
$N_{B}$ $=2k$ (12)
$n$ $=3k+1$ (13)
$L$ $=4k$ (14)
$C_{4}$ $=k$ (15)
$z_{\text{BipartiteFourCyclesNodePowerA}(\alpha)}$ $=k^{\alpha}+k$ (16)
$z_{\text{BipartiteFourCyclesNodePowerB}(\alpha)}$ $=2k$ (17)
Because $0<\alpha\leq 1$, the BipartiteFourCyclesNodePowerA statistic (16) will always be less than (or equal to, if $\alpha=1$ or $k=1$) the BipartiteFourCyclesNodePowerB statistic (17) for this family of graphs. The consequences for the interpretation of the corresponding parameters are discussed in Section 6.1.
5.1 Change statistics for the new statistics
The change statistic [40, 30, 25], that is, the difference in the statistic caused by adding a new edge $(i,j)$, for the four-cycles statistic (3) is
$\delta_{C4}(i,j)$ $=\sum_{k\in N(i)}L_{2}(j,k)$ (18)
$=\sum_{k\in N(j)}L_{2}(i,k)$ (19)
where $N(i)$ denotes the neighbours of node $i$, that is, nodes $k\neq i$ such that $x_{ik}=1$, or, equivalently, $d(i,k)=1$. The change statistic for the FourCyclesNodePower statistic (7) is then:
$\begin{split}\delta_{\text{FourCyclesNodePower}(\alpha)}(i,j)&=\left[C_{4}(i)+\delta_{C4}(i,j)\right]^{\alpha}-C_{4}(i)^{\alpha}\\ &+\left[C_{4}(j)+\delta_{C4}(i,j)\right]^{\alpha}-C_{4}(j)^{\alpha}\\ &+\sum_{k\in N(i)}\left[\left(C_{4}(k)+L_{2}(k,j)+x_{kj}L_{2}(k,i)\right)^{\alpha}-C_{4}(k)^{\alpha}\right]\\ &+\sum_{\left\{k\,:\,k\in N(j)\land\,k\not\in N(i)\right\}}\left[\left(C_{4}(k)+L_{2}(k,i)\right)^{\alpha}-C_{4}(k)^{\alpha}\right].\end{split}$ (20)
The four terms in equation (20) count the contributions from, respectively, node $i$, node $j$, the neighbours of node $i$, and the neighbours of node $j$ which are not also neighbours of node $i$. Note that in the third term (the contribution from neighbours of node $i$), a node $k$ can only be a neighbour of both node $i$ and node $j$ (that is, $k\in N(i)\land x_{kj}=1$) if the network is not bipartite.
The change statistic for the bipartite four-cycles statistic for the node set $A$ (8) is simpler than the general case (20), as we only count the contributions from the nodes in node set $A$. Specifically, we have:
$\begin{split}\delta_{\text{BipartiteFourCyclesNodePowerA}(\alpha)}(i,j)&=\left[C_{4}(i)+\delta_{C4}(i,j)\right]^{\alpha}-C_{4}(i)^{\alpha}\\ &+\sum_{k\in N(j)}\left[\left(C_{4}(k)+L_{2}(k,i)\right)^{\alpha}-C_{4}(k)^{\alpha}\right]\end{split}$ (21)
where $i\in A$, $j\in B$, and $k\in A$. The change statistic for BipartiteFourCyclesNodePowerB (9) is defined analogously.
5.2 Position of the new statistics in the dependence hierarchy
The configurations allowed in a model are determined by the assumptions as to which ties are allowed to depend on which other ties. Pattison & Snijders [22] (subsequently elucidated by Wang et al. [19] for bipartite networks, and more recently by Pattison et al. [77]) created a two-dimensional hierarchy of dependence assumptions, where the two dimensions are two facets of proximity: the form of the proximity condition, and the maximum distance between dependent ties. This two-dimensional hierarchy of dependence assumptions is illustrated in Fig. 3, showing a partial order structure, in which, if one dependence condition can be implied by another, it can be reached by a downwards path from the first to the second [19, p. 215].
Figure 3: Dependence hierarchy, adapted from Wang et al. [19, Fig. 3]. $\mathrm{SI}_{0}$ is not defined.
In order to describe the proximity conditions, it is useful to define some notation for neighbourhoods in a graph. In Section 5.1 we defined $N(i)$ as the neighbours of node $i$, that is, nodes $k\neq i$ such that $d(i,k)=1$. Following Pattison et al. [77], we now define $N_{q}(u)$, the $q$-neighbourhood of node $u$, as the set of nodes within geodesic distance $q$ of $u$, that is, nodes $v$ such that $d(u,v)\leq q$. Note that $N(u)$ as earlier defined in Section 5.1 is distinct from $N_{1}(u)$, the former being defined as nodes with a geodesic distance of exactly 1 from $u$ (and hence excluding $u$ itself), while the latter includes $u$ itself as $d(u,u)=0$. Further, we define $N_{q}(U)$, the $q$-neighbourhood of node set $U$, as $N_{q}(U)=\{v:v\in N_{q}(u)\text{ for some }u\in U\}$, that is, the set of nodes whose distance to some node in $U$ is no more than $q$.
The four forms of proximity conditions, describing the nature of the proximity between the neighbourhoods of two pairs of nodes whose respective tie variables are hypothesized to be conditionally dependent only if the proximity condition holds, can be summarised as follows [77], in order of increasing generality:
  1. 1.
    Strict $p$-inclusion. $\mathrm{SI}_{p}$ ($p\geq 1$) holds if the $p$-neighbourhood of each node in a pair includes both of the nodes in the other pair.
  2. 2.
    $p$-inclusion. $\mathrm{I}_{p}$ ($p\geq 0$) holds if the $p$-neighbourhood of each pair of nodes includes the other pair.
  3. 3.
    Partial $p$-inclusion. $\mathrm{PI}_{p}$ ($p\geq 0$) holds if the $p$-neighbourhood of one pair of nodes includes the other pair.
  4. 4.
    $p$-proximity. $\mathrm{D}_{p}$ ($p\geq 0$) holds if the $p$-neighbourhood of one pair of nodes has a non-empty intersection with the other pair.
The dependence condition $\mathrm{I}_{1}$ is equivalent to the widely-used “social circuit“ dependence assumption [19], in which two ties are conditionally dependent if they would form a four-cycle if both present [78, 79, 30, 37], and $\mathrm{PI}_{1}$ allows the “alternating pendant-triangle” statistics recently described by Pattison et al. [77].
The change statistic for FourCyclesNodePower (20) in computing the probability of a new edge $(i,j)$, depends only on edges between nodes in the two-neighbourhoods of $i$ and of $j$, since any two nodes in a four-cycle must be at a geodesic distance of at most two from each other (a four-cycle is a pair of nodes with two two-paths between them). Hence any edges on which the probability of edge $(i,j)$ depend must be in the two-neighbourhood of $\{i,j\}$. This puts the FourCyclesNodePower configuration in the dependence class two-proximity ($\mathrm{D}_{2}$): the $p$-neighbourhood (with $p=2$) of one pair of nodes has a non-empty intersection with the other pair [22, 77]. $\mathrm{D}_{2}$ is labelled with a red outline in the dependence hierarchy diagram shown in Fig. 3.
Because of the partial order structure of the dependence hierarchy, the stricter proximity forms for a fixed $p$ imply the more general ones (and $\mathrm{D}_{p}$ is the most general), and for a fixed proximity condition, smaller $p$ implies all the larger $p$ (so $\mathrm{D}_{1}$ implies $\mathrm{D}_{2}$ for instance). Hence, to show that the FourCyclesNodePower configuration, which is in $\mathrm{D}_{2}$, is not also in any more specific dependence class, it is sufficient to show that it is not in $\mathrm{D}_{1}$ and also not in $\mathrm{PI_{2}}$ (see Fig. 3). To do so, we can use Proposition 3 of Pattison et al. [77], which gives the properties that must hold for configurations implied by the dependence structures associated with each proximity condition.
Proposition 3(d) of Pattison et al. [77] states that “For $\mathrm{D}_{\mathrm{p}}$, each configuration is a subgraph in which every pair of edges lies on a path of length $\leq(p+2)$[77, p. 191]. So for a configuration to be in $\mathrm{D}_{1}$, every pair of edges must lie on a path of length three (or shorter). The FourCyclesNodePower configuration (Fig. 2) does not meet this requirement, since there are edges that do not lie on a path of length three or less: consider, for example, a pair of edges incident to the outermost node in the figure on two different four-cycles. These do not lie on a path of length three (but are on a path of length four, satisfying the requirement for $\mathrm{D}_{2}$ but not $\mathrm{D}_{1}$). Hence the FourCyclesNodePower configuration is not in dependence class $\mathrm{D}_{1}$.
Proposition 3(c) of Pattison et al. [77] states that “For $\mathrm{PI}_{\mathrm{p}}$, each configuration is a subgraph in which every pair of edges lies either on a cyclic walk of length $\leq(2p+2)$ or on a cyclic walk of length $\leq 2(p-r)+1$ with an additional path of length $\leq r+1$ attached to a node lying on the cyclic walk, for $10\leq r\leq p-$ 1[77, p. 191]. So, for $\mathrm{PI}_{2}$, each configuration is a subgraph in which every pair of edges is on a cyclic walk of length $\leq(2p+2)=6$, or on a cyclic walk of length $\leq 2(p-0)+1=5$ with an additional path of length $\leq 1$ attached to a node on the cyclic walk, or on a cyclic walk of length $\leq 2(p-1)+1=3$ with an additional path of length $\leq 2$ attached to a node on the cyclic walk. Again, we can see that the configuration for FourCyclesNodePower does not meet these conditions, considering a pair of maximally distant edges (those incident to the outermost nodes in two different four-cycles in Fig. 2). Such a pair of edges is neither on a six-cycle, and nor is it on a five-cycle with an additional path of length at most one or a three-cycle with an additional path of length at most two. Hence the FourCyclesNodePower configuration is not in dependence class $\mathrm{PI}_{2}$.
In the case of the BipartiteFourCyclesNodePower ($A$ and $B$) statistics for bipartite networks, the same reasoning applies (see Fig. 2).
5.3 Implementation
The new ERGM effects FourCyclesNodePower, BipartiteFourCyclesNodePowerA, and BipartiteFourCyclesNodePowerB are implemented in the EstimNetDirected [71] software, available from https://github.com/stivalaa/EstimNetDirected. BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB are also implemented, as b1np4c and b2np4c, as user-contrubuted statnet moodel terms [80, 81], available from https://github.com/stivalaa/ergm.terms.contrib. An example is described in Appendix C.
The b1nodematch and b2nodematch statistics [21] are also implemented in EstimNetDirected (as BipartiteNodematchAlphaA, BipartiteNodematchAlphaB, BipartiteNodematchBetaA, and BipartiteNodematchBetaB). These statistics are described in Section 7.2.
To count the number of unique type $A$ and $B$ nodes that are involved in four-cycles in a two-mode network, the CYPATH software (http://research.nii.ac.jp/~uno/code/cypath.html) [82] was used to enumerate all of the four-cycles (which are necessarily chordless in a bipartite network).
Scripts for data conversion, statistical analysis, and generating plots were written in R [83] using the igraph [84, 85] and ggplot2 [86] packages.
6 Simulation experiments
In order to investigate the effect of the BipartiteFourCyclesNodePowerA parameter, and compare it to that of the $\mathrm{K\text{-}C_{A}}$ [18] (known as XACA in MPNet and BipartiteAltKCyclesA in EstimNetDirected) parameter, we conducted some simulation experiments. In these experiments, bipartite networks with 750 nodes in node set $A$ and 250 nodes in set $B$ were simulated with the Edge, BipartiteAltStarsA [$\lambda=2$], and BipartiteAltStarsB [$\lambda=2$] parameters set to $-8.50$, $-0.20$, and $2.00$, respectively. In one set of experiments, the BipartiteAltKCyclesA parameter was varied from $-1.00$ to $1.00$ in increments of $0.01$, for each of three values of $\lambda$: $2$, $5$, and $10$. In another set of experiments, the BipartiteFourCyclesNodePowerA parameter was varied from $-1.00$ to $2.00$ in increments of $0.01$, for each of three values of $\alpha$: $1/10$, $1/5$, and $1/2$. The networks were simulated using the SimulateERGM program from the EstimNetDirected software package, using the tie/no-tie (TNT) sampler [87], with a burn-in of $10^{7}$ iterations and an interval of $10^{5}$ iterations between each of 100 samples, to ensure that samples are drawn from the equilibrium ERGM distribution, and are not too autocorrelated.
The results of these simulations are shown in Fig. 4. Using the BipartiteAltKCyclesA parameter (left column) results in phase transition or near-degeneracy behaviour, with the statistic showing a sudden sharp increase at a critical value of the parameter. At this critical value, the graph density (top left plot) also sharply increases, as does the number of four-cycles (bottom left plot), which, at parameter values less than the critical value, hardly increased at all. This behaviour is similar to that of the simple Markov (edge-triangle) model described in Koskinen & Daraganova [37], and is characteristic of near-degeneracy in ERGMs. This can prevent estimation of models which contain parameters that cause this behaviour, and yet occurs in this case even when using an “alternating” statistic, designed to try to avoid such behaviour [18]. Note that changing the $\lambda$ parameter appears merely to change the maximum value of the statistic; it does not remove or “smooth out” the phase transition. In contrast, when using the new BipartiteFourCyclesNodePowerA parameter (right column) the phase transition behaviour is less apparent even for the highest value of $\alpha$ ($1/2$), and can be smoothed out further as we decrease $\alpha$. So, by appropriately setting $\alpha$, we can use the BipartiteFourCyclesNodePowerA parameter to generate smoothly varying numbers of four-cycles, without suddenly tipping from a low-density low-clustering regime to a high-density high-clustering regime with nothing in between, which did not seem to be possible on this example with the BipartiteAltKCyclesA parameter.
Figure 4: Effect of varying the BipartiteAltKCyclesA parameter (left) and BipartiteFourCyclesNodePowerA parameter (right) on the Edge statistic (top), the statistic corresponding to the parameter itself (middle) and the FourCycles statistic (bottom). Each graph plots the mean value of the statistic (over 100 simulated networks) for three different values of the $\lambda$ or $\alpha$ parameter for BipartiteAltKCyclesA and BipartiteFourCyclesNodePowerA, respectively.
6.1 Interpretation of the new parameters
Interpretation of the FourCyclesNodePower parameter is that a positive value increases the number of four-cycles and a negative value decreases the number of four-cycles, relative to a value of zero. A smaller value of the exponent $\alpha$ means that additional four-cycles including the same node contribute less than if those cycles involved distinct nodes.
Interpretation of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters is rather more complicated and is illustrated in Fig. 5 and Fig. 6. Note that in a bipartite network, any four-cycle must contain two nodes in node set $A$ and two nodes and in node set $B$. So how can we get more four-cycles in one mode than the other? The answer is that the four-cycle counts for the two must be equal, but the weighted node-oriented four-cycle counts (8) and (9) can differ. As discussed in Section 5, the statistic BipartiteFourCyclesNodePowerA (8) is maximised by having four-cycles involving distinct pairs of nodes in node set $A$ (rather than many four-cycles involving the same node in node set $A$). If the BipartiteFourCyclesNodePower parameter for $A$ is positive and for $B$ is zero (or negative) then we tend to get more mode $A$ nodes involved in four-cycles, with the same mode $B$ nodes participating in many four-cycles, since the statistic is higher by having different nodes in the four-cycles, than for having the same node involved in many four-cycles. In the examples illustrated in Fig. 6, this results in the mode $A$ nodes being part of a denser core with lots of four-cycles with a smaller number of $B$ nodes, resulting in isolated $B$ nodes. And vice versa for $A$ zero (or negative) and $B$ positive (“zero.pos” and “neg.pos”; these are perhaps clearer as there are more $A$ nodes than $B$ nodes in the network). Particularly in the “neg.pos” case, we can see a core of mode $B$ nodes connected to a small number of central mode $A$ nodes in four-cycles (as well as others not involved in four-cycles) and many mode $A$ isolates. If BipartiteFourCyclesNodePower for both $A$ and $B$ are positive then there are even more four-cycles, but they are more evenly distributed between the $A$ and $B$ nodes.
Equation (10) implies that there are two degrees of freedom for the three parameters; for example we can include both BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB in a model, but not also FourCyclesNodePower since it is the sum of the other two.
Figure 5: Effect of different combinations of negative, zero, and positive values of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters on (top to bottom) the BipartiteFourCyclesNodePowerA statistic, the BipartiteFourCyclesNodePowerB statistic, and the FourCycles statistic. Box plots show the statistics of 100 simulated networks. The simulated bipartite networks have 100 nodes in node set A and 50 nodes in node set B and are simulated with common parameters $(\mathrm{Edge},\mathrm{BipartiteAltStarsA}~{}[\lambda=2],\mathrm{BipartiteAltStarsB}~{}[\lambda=2])=(-6.0,-0.4,1.0)$. For $\mathrm{BipartiteFourCyclesNodePowerA}~{}[\alpha=1/5]$ and $\mathrm{BipartiteFourCyclesNodePowerB}~{}[\alpha=1/5]$ the negative (“neg”) parameter value is $-1.5$ and the positive (“pos”) parameter value is $6.5$.
Figure 6: Examples of networks simulated with different combinations of negative, zero, and positive values of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters, drawn from the simulations shown in Fig. 5. Nodes in node set A are shown as red circles, and nodes in node set B are shown as blue squares.
7 Empirical applications
In this section we demonstrate the application of the new statistics in ERGM models of a selection of empirical two-mode networks ranging in size from hundreds of nodes to hundreds of thousands of nodes. Network summary statistics of the empirical networks are shown in Table 4 (four-cycle counts are shown in Table E1). The largest of these networks is larger than any of the networks in the literature survey (Table LABEL:tab:literature).
Table 4: Bipartite graph summary statistics. “Bipartite c.c” is the Robins-Alexander bipartite clustering coefficient [9].
Network $\mathrm{N_{A}}$ $\mathrm{N_{B}}$ Mean degree Bipartite
A B c.c.
Inouye-Pyke pollinators 91 42 3.09 6.69 0.20856
St Louis crime 870 557 1.71 2.67 0.05041
Norwegian director interlock 1 495 367 1.23 5.00 0.09651
Robertson pollinators 1 428 456 10.68 33.45 0.10876
Australian director interlock 9 971 2 087 1.35 6.45 0.08438
Scientific collaborations 16 726 22 015 3.50 2.66 0.22822
International director interlock 321 869 34 769 1.17 10.84 0.12617
Three of the empirical network examples are director interlock networks [88], where nodes represent directors and companies, with an edge indicating that a director sits on a company board. Two of the empirical network examples are plant-pollinator webs, where edges represent an interaction between a plant and a pollinator, one network represents participation of actors in crime events, and one network represents scientific collaboration (co-authorship of scientific papers).
All of the ERGM models in this section were estimated with EstimNetDirected [71] using the simplified EE algorithm [28] with the improved fixed density (IFD) sampler [26] with 64 parallel estimation runs. Only models in which all runs converge and for which the model is non-degenerate (the observed statistics for effects in the model are with the 95% region of statistics simulated from the model) are presented. Simulations from the estimated models, used for the degeneracy check and goodness-of-fit testing, were done with the SimulateERGM program from the EstimNetDirected software package, with the tie/no-tie (TNT) sampler [87], with the exception of the Inouye-Pyke pollinator web, in which the IFD sampler was used.
7.1 Inouye-Pyke pollinator web
The Inouye-Pyke pollinator web [89] data was downloaded from the Interaction Web Database (IWDB) at https://www.nceas.ucsb.edu/interactionweb/html/inouye_1988.html (accessed 5 June 2018). (Note that the IWDB has since relocated and this data is now available from http://www.ecologia.ib.usp.br/iwdb/html/inouye_1988.html.) In this network, the node set $A$ represents pollinators and node set $B$ represents plants, and an edge represents an interaction between a pollinator and a plant.
Table 5 shows two ERGM models of Inouye-Pyke pollinator web. Model 1 includes the BipartiteAltKCyclesB term (models with BipartiteAltKCyclesA or FourCycles did not converge). Model 2 includes the new BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB terms. The former is found to be negative (but not statistically significant) while the latter is positive and statistically significant. This indicates a statistically significant tendency towards having multiple four-cycles involving distinct pairs of plants (node set $B$), and together with the negative or zero (not significant) BipartiteFourCyclesNodePowerA parameter, it models a network with pollinators (node set $A$) participating in multiple four-cycles. This is similar to the “zero.pos” or “neg.pos” case in Fig. 5 and Fig. 6, not coincidentally, since that simulation was based on this model.
Table 5: ERGM parameter estimates for the Inouye-Pyke pollinator web with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”.
Effect Model 1 Model 2
Edge $\mathbf{\underset{(-6.366,-5.598)}{-5.982}}$ $\mathbf{\underset{(-6.470,-5.989)}{-6.229}}$
BpAltStarsA [$\lambda=2$] $\underset{(-0.552,2.397)}{0.923}$ $\underset{(-1.674,0.878)}{-0.398}$
BpAltStarsB [$\lambda=5$] $\underset{(-0.493,1.597)}{0.552}$ $\mathbf{\underset{(0.319,1.706)}{1.013}}$
BpAltKCyclesB [$\lambda=2$] $\underset{(-0.322,0.332)}{0.005}$
BpFourCyclesNodePowerA [$\alpha=1/5$] $\underset{(-4.443,1.662)}{-1.391}$
BpFourCyclesNodePowerB [$\alpha=1/5$] $\mathbf{\underset{(1.670,11.401)}{6.535}}$
The positive and statistically significant BipartiteAltStarsB parameter indicates a tendency for centralisation on plant nodes, that is, plants that attract many pollinators.
The fit of the two models to the cycle length distribution of the network is shown in Fig. 7(a); Model 2 (right) has a slightly better fit to the distribution (and in particular four-cycles) than Model 1 (left). Fig. D1 shows more extensive goodness-of-fit plots for the two models. Model 2 has a better fit to all of the statistics included in the goodness-of-fit plots: degree distributions, giant component size, geodesic distance distribution, dyadwise shared partner distribution, and the bipartite clustering coefficients, as well as the cycle length distribution.
Figure 7: Cycle length distribution goodness-of-fit plots for the empirical network examples. Plots in the left column are from models using the BipartiteAltKCycles parameter(s), where possible (no such model could be found for the St Louis crime network, so a simple model with no four-cycles effects is used), and plots in the right column are from models using the new BipartiteFourCyclesNodePower parameter(s). Observed network statistics are plotted as red points (joined by red lines as a visual aid) with box plots of the statistics of 100 simulated networks.
We also estimated Model 1 using MPNet (Table E2), and found the same parameter signs, but with smaller standard errors, so that XACB (BipartiteAltKCyclesB) is positive and significant, as are both XASA and XASB (BipartiteAltStarsA and BipartiteAltStarsB). This would appear to be a case of the EE algorithm (used in EstimNetDirected) having lower statistical power than the stochastic approximation algorithm [24] (used in MPNet) if an insufficient number of estimations runs are used [90, 91]. A positive and statistically significant estimate of the XACB parameter indicates a tendency for multiple pollinators to share the same plants (even in addition to the positive XASB (BipartiteAltStarsB) parameter indicating the presence of individual plant species that attract many pollinators).
7.2 St Louis crime network
The St Louis crime network data [92] was obtained from the networkdata R package [93]. In this data, the nodes in node set $A$ represent people, and the nodes in node set $B$ represent crime events. In the original data, an edge can represent participation as a victim, suspect, witness or dual (both victim and suspect). Here, we consider any of these edge types to represent participation in the crime event. Individuals also have their gender recorded.
Four models of this network are shown in Table 6. Model 1 is a “baseline” model, with only edge, isolates, and star (degree distribution) effects. We were unable to find any converged models using the FourCycles or BipartiteAltKCycles (A or B) effects. Model 2 adds the BipartiteFourCyclesNodePower (A and B) effects. Fig. 7(b) shows the fit of Model 1 (left) and Model 2 (right) to the cycle length distribution of this network. Model 1 does not fit four-cycles well (but fits six-cycles, eight-cycles, and ten-cycles relatively well), while Model 2 fits four-cycles and six-cycles well, with the fit becoming progressively worse for longer cycle lengths (as it does for Model 1 with cycle lengths greater than ten).
Table 6: ERGM parameter estimates for the St Louis crime network with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”
Effect Model 1 Model 2 Model 3 Model 4
Edge $\mathbf{\underset{(-12.589,-12.102)}{-12.345}}$ $\mathbf{\underset{(-11.825,-11.349)}{-11.587}}$ $\mathbf{\underset{(-11.685,-11.193)}{-11.439}}$ $\mathbf{\underset{(-11.226,-10.723)}{-10.974}}$
BpIsolatesA $\mathbf{\underset{(-10.449,-4.600)}{-7.524}}$ $\mathbf{\underset{(-10.276,-3.948)}{-7.112}}$ $\mathbf{\underset{(-10.335,-3.885)}{-7.110}}$ $\mathbf{\underset{(-10.420,-3.809)}{-7.114}}$
BpIsolatesB $\underset{(-10.860,2.668)}{-4.096}$ $\underset{(-10.767,2.857)}{-3.955}$ $\underset{(-10.966,3.306)}{-3.830}$ $\underset{(-10.700,3.710)}{-3.495}$
BpAltStarsA [$\lambda=2$] $\mathbf{\underset{(2.042,4.592)}{3.317}}$ $\mathbf{\underset{(1.443,4.259)}{2.851}}$ $\mathbf{\underset{(1.382,4.314)}{2.848}}$ $\mathbf{\underset{(1.362,4.339)}{2.850}}$
BpAltStarsB [$\lambda=2$] $\underset{(-0.709,2.187)}{0.739}$ $\underset{(-0.981,1.980)}{0.499}$ $\underset{(-1.004,2.046)}{0.521}$ $\underset{(-1.147,2.048)}{0.450}$
BpFourCyclesNodePowerA [$\alpha=1/5$] $\underset{(-2.615,3.506)}{0.446}$ $\underset{(-2.664,3.558)}{0.447}$ $\underset{(-2.709,3.597)}{0.444}$
BpFourCyclesNodePowerB [$\alpha=1/3$] $\underset{(-0.582,3.601)}{1.510}$ $\underset{(-0.625,3.641)}{1.508}$ $\underset{(-0.659,3.689)}{1.515}$
BpActivityA female $\underset{(-0.758,0.608)}{-0.075}$ $\underset{(-2.385,1.245)}{-0.570}$
BpNodematchA gender [$\beta=0.1$] $\underset{(-2.028,1.530)}{-0.249}$
BpNodematchA female [$\beta=0.1$] $\underset{(-1.580,1.707)}{0.063}$
BpNodematchA male [$\beta=0.1$] $\underset{(-3.435,1.883)}{-0.776}$
More extensive goodness-of-fit plots for the two models are shown in Fig. D2. Model 2 fits well on all but geodesic distance distribution and giant component size. Notably it has better fit on the dyadwise shared partner distribution and the bipartite clustering coefficients than does Model 1, as a consequence of the former’s better fit to four-cycles.
Model 3 adds terms for gender homophily (including the term “BipartiteActivityA female” to capture a tendency for female actors to be more or less active in the network), while Model 4 adds terms for differential homophily on gender, that is, different homophily terms for male and female actors. BipartiteNodematchA [$\beta$] and BipartiteNodematchB [$\beta$] are the same as the statnet terms b1nodematch($\beta$) and b2nodematch($\beta$), as defined in Bomiriya et al. [21], for modelling homophily from an edge-centred view, in bipartite networks. The b1nodematch($\beta$) term is described by the ergm R package documentation as “half the sum of the number of two-paths with two first-mode nodes with that attribute as the two ends of the two path raised to the power beta for each edge in the network” [48, ’b1nodematch-ergmTerm’], in other words, equation (7) of Bomiriya et al. [21]. This homophily statistic is edge-centred, as for every edge in a bipartite network, it counts the number of matching two-stars containing the edge (that is, pairs of nodes in one mode with the same attribute value that are connected to a common node in the other mode via a two-path containing that edge). This count for each such pair of nodes is then raised to the power $\beta$, and the sum is divided by two since each two-path contains two edges [21, p. 6]. Using these terms instead of measuring homophily in a more straightforward way by simply counting the number of two-paths between nodes with the same value of an attribute is advantageous for the two reasons discussed in Bomiriya et al. [21, p. 5]. First, the diminishing returns of additional two-paths connecting a pair of nodes with the same attribute that are already connected by a two-path means the b1nodematch($\beta$) term may be more appropriate from a modelling perspective, and, second, the simple two-path counting method can lead to problems of model degeneracy, which is alleviated by the use of an appropriate exponent $\beta$ in the b1nodematch($\beta$) term.
The only effects that are found to be statistically significant (other than Edge) are BipartiteIsolatesA and BipartiteAltStarsA. The negative value of the former is likely a reflection of how the network was constructed: isolated people nodes are unlikely to occur as the network was constructed by a snowball sampling process from an initial set of crime reports. The positive value of the latter confirms a centralisation effect on person nodes, that is, there are people who are involved in a disproportionately large number of crime events. (There is no significant such effect for crime nodes, that is, crimes involving a disproportionately large number of people). None of the gender-related parameters are statistically significant, so we cannot draw any conclusions from this model as to the influence of gender in this network.
7.3 Robertson pollinator web
The Robertson pollinator network [94, 95] data was downloaded from the Interaction Web Database (IWDB) at http://www.ecologia.ib.usp.br/iwdb/html/robertson_1929.html (accessed 19 April 2022). The node set $A$ represents pollinators and the node set $B$ represents plants, with an edge representing a plant-pollinator interaction. We were unable to find any converged model for this network, even with only edge and alternating-$k$-star terms. One reason that this network could be problematic for ERGM modelling is that it has at least an order of magnitude higher bipartite density than the other networks (with the exception of the Inouye-Pyke pollinator web, which has only 133 nodes, while this network has 1 884 nodes); this is also evident in the order of magnitude higher mean degrees for the Robertson pollinator web (Table 4). ERGMs are typically applied to social networks which are (usually) sparse, especially in the case of large networks [71], and there can be difficulties in trying to estimate ERGM parameters if the network is too dense; for example the neural networks for which ERGM models could not be estimated in Stivala & Lomi [96] and Stivala [97] also have density and mean degree at least an order of magnitude larger than the other biological networks [97, Table 1], for which ERGMs could be estimated. In those cases, the problem was overcome by using either the “Tapered ERGM” [31, 35, 98] or the LOLOG model [60, 59].
7.4 Norwegian director interlock network
The Norwegian director interlock network data [99] was downloaded from the supporting website described in the paper (accessed 29 June 2020). Note that this website is apparently no longer available, but we have retained a copy of the data which we downloaded. In this data, the nodes in node set $A$ represent company directors, and the nodes in node set $B$ represent companies; an edge represents a director sitting on the board of a company. We use only the network for 1 August, 2009, as also done in Opsahl [10]. There is also an attribute for the gender of directors, and location data (postcode and city) for companies.
Table 7 shows four models of this network. Model 1 is a structural model using the BipartiteAltKCyclesB term (models with BipartiteAltKCyclesA did not converge), while Model 2 is a structural model using the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB terms. Model 3 and Model 4 add terms for geographical and gender homophily, with Model 4 testing for differential homophily on gender. These models use the BipartiteNodematchA [$\beta$] and BipartiteNodematchB [$\beta$] terms, as discussed in Section 7.2. The only statistically significant parameters (apart from Edge) are BipartiteAltStarsA and BipartiteFourCyclesNodePowerA (although the latter is not significant in Model 4, where terms for differential gender homophily are introduced). The negative BipartiteAltStarsA parameter indicates a lack of centralisation on the director degree distribution. The positive BipartiteFourCyclesNodePowerA (and negative, but not statistically significant, BipartiteFourCyclesNodePowerB) parameter means that there is a tendency towards having more directors than boards involved in four-cycles, with the same company boards participating in multiple four-cycles.
Table 7: ERGM parameter estimates for the Norwegian director interlock network with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”
Effect Model 1 Model 2 Model 3 Model 4
Edge $\mathbf{\underset{(1.950,2.451)}{2.200}}$ $\mathbf{\underset{(2.675,3.112)}{2.893}}$ $\mathbf{\underset{(2.487,2.930)}{2.709}}$ $\mathbf{\underset{(1.153,1.599)}{1.376}}$
BpAltStarsA [$\lambda=2$] $\mathbf{\underset{(-4.927,-1.758)}{-3.342}}$ $\mathbf{\underset{(-5.320,-1.995)}{-3.658}}$ $\mathbf{\underset{(-5.655,-2.087)}{-3.871}}$ $\mathbf{\underset{(-5.675,-2.070)}{-3.872}}$
BpAltStarsB [$\lambda=2$] $\underset{(-9.662,4.167)}{-2.748}$ $\underset{(-7.752,1.551)}{-3.101}$ $\underset{(-8.114,1.751)}{-3.181}$ $\underset{(-7.986,1.712)}{-3.137}$
BpAltKCyclesB [$\lambda=5$] $\underset{(-0.467,0.470)}{0.001}$
BpFourCyclesNodePowerA [$\alpha=1/3$] $\mathbf{\underset{(0.660,4.364)}{2.512}}$ $\mathbf{\underset{(0.070,4.693)}{2.382}}$ $\underset{(-0.155,4.902)}{2.373}$
BpFourCyclesNodePowerB [$\alpha=1/3$] $\underset{(-2.118,1.626)}{-0.246}$ $\underset{(-2.603,2.347)}{-0.128}$ $\underset{(-2.934,2.699)}{-0.118}$
BpActivityA female $\underset{(-0.796,1.747)}{0.476}$ $\underset{(-1.265,5.625)}{2.180}$
BpNodematchA gender [$\beta=0.1$] $\underset{(-2.511,3.085)}{0.287}$
BpNodematchA female [$\beta=0.1$] $\underset{(-2.331,1.514)}{-0.409}$
BpNodematchA male [$\beta=0.1$] $\underset{(-3.308,7.763)}{2.228}$
BpNodematchB city [$\beta=0.1$] $\underset{(-0.783,1.977)}{0.597}$ $\underset{(-0.835,2.029)}{0.597}$
The fit of Model 1 (left) and Model 2 (right) to the cycle length distribution of the network is shown in Fig. 7(c). Neither fit the distribution well, but Model 2 is far closer to fitting the number of four-cycles well than Model 1 is.
Further goodness-of-fit plots are shown in Fig. D3. Model 1 fits the degree distributions well, but is not good on giant component size, geodesic distance distribution, dyadwise shared partner distribution, and, as already described, cycle length distribution (and hence bipartite clustering coefficients). Model 2 improves on Model 1 by having far better fit on the dyadwise shared partner distribution and (as already discussed), the number of four-cycles, and hence also the Robins-Alexander (but not Opsahl) bipartite clustering coefficient.
7.5 Australian director interlock network
Australian company director data was obtained from the Thomson Reuters Connect 4 Boardroom database https://www.thomsonreuters.com.au/en-au/products/connect4.html (accessed 14 September 2022 via Swinburne University of Technology institutional subscription). This data includes information on the country, gender and age of directors, and which companies they are directors of. This data was joined with other data from the Australian Securities Exchange (ASX), specifically the company directory https://www.asx.com.au/asx/research/ASXListedCompanies.csv (accessed 5 October 2022) and the “foreign entity report” for September 2022 https://www.asx.com.au/listings/listing-considerations/listing-requirements/foreign-entity-data (accessed 23 October 2022). From these data sources, a director interlock network was constructed, with nodes in mode $A$ representing directors, and nodes in mode $B$ representing companies. This network is the director interlock network for all companies listed on the ASX (as of September 2022), and also contains node attributes for the gender, age, and country of residence of directors, and the industry group (GICS), listing date, market capitalisation, and foreign country incorporation of companies. This is the same data used in earlier work [100], where more details, including descriptive statistics and simpler ERGM models, can be found.
Table 8 shows three ERGM models of the Australian director interlock network. Model 1 and Model 2 include the BipartiteAltKCyclesB parameter (models containing BipartiteAltKCyclesA did not converge), while Model 3 includes the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters. Just as for the Norwegian director interlock network (Section 7.4), BipartiteAltStarsA is negative and significant in all models, BipartiteAltKCyclesB is not significant in models that contain it (Models 1 and 2), and BipartiteFourCyclesNodePowerA is positive and significant with BipartiteFourCyclesNodePowerB not significant in Model 3. Hence, just as for the Norwegian director interlock network, there is a tendency towards having more directors than boards involved in four-cycles, with the same company boards participating in multiple four-cycles.
Table 8: ERGM parameter estimates for the Australian director interlock network with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”
Effect Model 1 Model 2 Model 3
Edge $\mathbf{\underset{(-3.603,-3.385)}{-3.494}}$ $\mathbf{\underset{(-2.369,-2.083)}{-2.226}}$ $\mathbf{\underset{(-1.634,-1.339)}{-1.486}}$
BpAltStarsA [$\lambda=1.1$] $\mathbf{\underset{(-7.255,-2.991)}{-5.123}}$ $\mathbf{\underset{(-7.952,-2.295)}{-5.124}}$ $\mathbf{\underset{(-8.549,-2.528)}{-5.538}}$
BpAltStarsB [$\lambda=5$] $\underset{(-1.806,0.838)}{-0.484}$ $\underset{(-2.281,1.248)}{-0.517}$ $\underset{(-1.658,0.044)}{-0.807}$
BpAltKCyclesB [$\lambda=5$] $\underset{(-0.459,0.286)}{-0.086}$ $\underset{(-0.577,0.436)}{-0.071}$
BpFourCyclesNodePowerA [$\alpha=1/4$] $\mathbf{\underset{(0.194,4.206)}{2.200}}$
BpFourCyclesNodePowerB [$\alpha=1/4$] $\underset{(-1.448,2.939)}{0.746}$
BpActivityA female $\underset{(-0.209,1.524)}{0.657}$ $\underset{(-1.788,0.507)}{-0.640}$ $\underset{(-2.005,0.323)}{-0.841}$
BpContinuousActivityA age $\mathbf{\underset{(0.001,0.020)}{0.010}}$ $\underset{(-0.002,0.023)}{0.010}$ $\underset{(-0.003,0.024)}{0.010}$
BpActivityA country.Australia $\mathbf{\underset{(-1.627,-0.884)}{-1.256}}$ $\mathbf{\underset{(-2.001,-0.516)}{-1.258}}$ $\mathbf{\underset{(-1.878,-0.117)}{-0.998}}$
BpActivityB country.Australia $\underset{(-2.104,0.417)}{-0.844}$ $\underset{(-2.573,0.889)}{-0.842}$ $\underset{(-2.528,0.858)}{-0.835}$
BpActivityB industryGroup.Banks $\underset{(-1.615,2.661)}{0.523}$ $\underset{(-2.558,3.591)}{0.516}$ $\underset{(-2.339,3.212)}{0.436}$
BpActivityB industryGroup.Materials $\underset{(-1.217,0.099)}{-0.559}$ $\underset{(-1.440,0.322)}{-0.559}$ $\underset{(-1.495,0.479)}{-0.508}$
BinaryPairInteraction gender.F country.Australia $\underset{(-0.533,0.535)}{0.001}$ $\underset{(-0.687,0.708)}{0.011}$ $\underset{(-0.561,0.616)}{0.028}$
BinaryPairInteraction gender.F industryGroup.Banks $\underset{(-2.042,2.985)}{0.472}$ $\underset{(-3.185,4.083)}{0.449}$ $\underset{(-2.748,3.563)}{0.408}$
BinaryPairInteraction gender.F industryGroup.Materials $\underset{(-1.478,0.535)}{-0.471}$ $\underset{(-1.773,0.855)}{-0.459}$ $\underset{(-1.820,0.939)}{-0.441}$
Matching country $\mathbf{\underset{(0.804,3.523)}{2.163}}$ $\mathbf{\underset{(0.280,4.043)}{2.162}}$ $\mathbf{\underset{(0.089,4.113)}{2.101}}$
BpContinuousActivityB ListingYear $\underset{(-0.001,0.001)}{-0.000}$ $\underset{(-0.002,0.002)}{-0.000}$ $\underset{(-0.002,0.002)}{-0.000}$
BpContinuousActivityB logMarketCap $\underset{(-0.011,0.154)}{0.072}$ $\underset{(-0.037,0.179)}{0.071}$ $\underset{(-0.048,0.193)}{0.072}$
BpNodematchA country [$\beta=0.1$] $\mathbf{\underset{(0.827,6.152)}{3.489}}$ $\mathbf{\underset{(0.067,6.921)}{3.494}}$ $\underset{(-0.615,7.390)}{3.388}$
BpNodematchB country [$\beta=0.1$] $\mathbf{\underset{(-2.900,-0.204)}{-1.552}}$ $\underset{(-3.358,0.256)}{-1.551}$ $\underset{(-3.375,0.423)}{-1.476}$
BpNodematchB industryGroup [$\beta=0.1$] $\mathbf{\underset{(1.018,2.443)}{1.731}}$ $\mathbf{\underset{(0.794,2.668)}{1.731}}$ $\mathbf{\underset{(0.267,2.550)}{1.409}}$
BpNodematchA gender [$\beta=0.1$] $\underset{(-1.139,2.193)}{0.527}$
BpNodematchA female [$\beta=0.1$] $\underset{(-1.562,2.756)}{0.597}$ $\underset{(-1.689,2.843)}{0.577}$
BpNodematchA male [$\beta=0.1$] $\underset{(-4.166,1.417)}{-1.374}$ $\underset{(-4.502,1.247)}{-1.627}$
Figure 7(d) shows the fit of Model 2 (left) and Model 3 (right) to the cycle length distribution. Neither fit the distribution well, but Model 3 (with the new BipartiteFourCyclesNodePower parameters) has a far better fit to the number of four-cycles than Model 2 (with BipartiteAltKCyclesB) does.
Further goodness-of-fit plots are shown in Fig. D4. Both models have reasonable fit to the degree and geodesic distance distributions. Model 3 (with the new BipartiteFourCyclesNodePower parameters) has a better fit to the dyadwise shared partner distribution, and, as a consequence of fitting four-cycles better than Model 2, a better (but still not good) fit on the Robins-Alexander bipartite clustering coefficient (but not the Opsahl clustering coefficient, which depends on fitting six-cycles well).
These models also reproduce the attribute effects found in the simpler ERGM models of this network described in Stivala et al. [100]. Older directors tend to be on more boards (positive and significant BipartiteContinuousActivityA age, although this is only significant in Model 1), directors resident in Australia tend to be on fewer boards than those that are not (negative and significant BipartiteActivityA country.Australia), directors tend to be resident in the same country as the country of incorporation of companies whose boards they sit on (positive and significant Matching country), and directors on a board tend to be from the same country (positive and significant BipartiteNodematchA country in Models 1 and 2).
Although the models in Stivala et al. [100] found that larger market capitalisation is associated with larger boards (positive and significant BipartiteContinuousActivityB logMarketCap) and that foreign incorporated boards tend to have more directors, the equivalent effects here are not significant (BipartiteContinuousActivityB logMarketCap is positive but not significant and BipartiteActivityB country.Australia is negative but not significant).
In Stivala et al. [100], the simple homophily term BipartiteTwoPathMatchingA was used to test for gender homophily, but none of the gender effects were statistically significant. Here we use instead the BipartiteNodematchA [$\beta$] term, as discussed in Section 7.2, to test for gender homophily (Model 1) and differential gender homophily (Models 2 and 3). Again, however, none of the gender effects are statistically significant.
Models with the simple BipartiteTwoPathMatchingB term to test for country and industry homophily on boards (that is, to test if boards in the same country or industry tend to share directors) did not converge, but this problem was solved by using the BipartiteNodematchB [$\beta$] term instead. The effect for industry group homophily is positive and significant in all models, confirming that companies in the same industry group tend to be connected by shared directors. The effect for country homophily is negative (and statistically significant in Model 1 only).
7.6 Scientific collaborations
The scientific collaboration data is a co-authorship network of preprints in the Condensed Matter section of the arXiv e-print archive [101, 102, 103]. The nodes in node set $A$ represent authors, and the nodes in node set $B$ represent papers, with an edge between two nodes representing authorship on a paper. The data was downloaded from https://toreopsahl.com/datasets/#newman2001 (accessed 17 April 2024); note that we use the two-mode (bipartite) version of this network as used in Opsahl [10], rather than the one-mode projection onto authors as is frequently used.
Table 9 shows two ERGM models for this network. Model 1 is a baseline model with only Edge, IsolateEdges, and degree distribution (alternating $k$-star) parameters. Model 2 adds the BipartiteAltKCyclesB parameter. Models including the FourCycles or BipartiteAltKCyclesA parameters did not converge, and nor did models including the new FourCyclesNodePower, BipartiteFourCyclesNodePowerA or BipartiteFourCyclesNodePowerB parameters.
Table 9: ERGM parameter estimates for the scientific collaboration network with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”.
Effect Model 1 Model 2
Edge $\mathbf{\underset{(-9.873,-8.785)}{-9.329}}$ $\mathbf{\underset{(-10.752,-10.172)}{-10.462}}$
IsolateEdges $\mathbf{\underset{(0.110,1.956)}{1.033}}$ $\mathbf{\underset{(0.438,2.323)}{1.380}}$
BpAltStarsA [$\lambda=2$] $\mathbf{\underset{(0.821,1.234)}{1.028}}$ $\mathbf{\underset{(0.820,1.236)}{1.028}}$
BpAltStarsB [$\lambda=2$] $\mathbf{\underset{(-1.003,-0.412)}{-0.708}}$ $\mathbf{\underset{(0.830,2.118)}{1.474}}$
BpAltKCyclesB [$\lambda=5$] $\mathbf{\underset{(-1.144,-0.611)}{-0.878}}$
The positive and significant IsolateEdges parameter reflects the presence of more single-author papers than expected, given the other effects in the model, which include terms for the numbers of authors per paper and numbers of papers per author.
The negative and significant BipartiteAltKCyclesB parameter (in conjunction with the positive and significant BipartiteAltStarsB parameter reflecting the existence of papers with many authors) indicates that there is a tendency against multiple authors being on multiple papers together. Similar to the situation described for the top 500 Australian listed companies director interlock network described by Wang et al. [18, pp. 23–24], this, together with the presence of authors with many papers (positive and significant BipartiteAltStarsA parameter) suggests that there are highly active authors who tend to be on papers that do not otherwise share authors.
The goodness-of-fit plots for these models (Fig. D5) show that they have reasonable fit on mode $B$ (paper) degree, but poor on mode $A$ (author) degree (for low degree). The models also do not fit well to the dyadwise shared partner distribution (they do not reproduce the long tail), cycle length distribution, or the bipartite clustering coefficients. Model 2, despite including the BipartiteAltKCyclesB parameter, does not fit the dyadwise shared partner or four-cycles distributions (and hence bipartite clustering coefficients) any better than the baseline Model 1.
This network is therefore an example where the bipartite alternating $k$-cycle parameters do not improve the fit to four-cycles, and neither do the new BipartiteFourCyclesNodePower parameters, as converged models with the latter could not be found.
7.7 International director interlock network
The international director interlock network described in Evtushenko & Gastner [104] was downloaded from [105]. In this network, nodes in node set $A$ represent directors, nodes in node set $B$ represent companies, and an edge represents a director sitting on the board of a company. Director nodes have attributes for age and gender, while company nodes have attributes for country, sector, industry, and number of employees. This data was also used in earlier work [100], in which the ERGM model shown as Model 1 in Table 10 was first presented.
Table 10 shows three models of this network. Model 1 reproduces the model described in Stivala et al. [100], but estimated with 64 parallel estimation runs, rather than the 20 used in Stivala et al. [100]. Model 2 adds homophily effects for country, industry and sector on companies, and gender on directors, while Model 3 further adds effects for differential gender homophily. In Stivala et al. [100], none of these effects could be included as models including any of the simple homophily parameters BipartiteTwoPathMatchingA and BipartiteTwoPathMatchingB would not converge. By using the BipartiteNodematchB [$\beta$] and BipartiteNodematchA [$\beta$] parameters instead (see Section 7.2), however, these homophily effects can be included in the model. None of the models contain any of the FourCycles, BipartiteAltKCycles or BipartiteFourCyclesNodePower effects, as models containing any of these parameters did not converge.
Table 10: ERGM parameter estimates for the international director interlock network with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”
Effect Model 1 Model 2 Model 3
Edge $\mathbf{\underset{(-10.702,-10.080)}{-10.391}}$ $\mathbf{\underset{(-10.248,-9.710)}{-9.979}}$ $\mathbf{\underset{(-10.159,-9.623)}{-9.891}}$
IsolateEdges $\underset{(-0.380,0.205)}{-0.087}$ $\underset{(-0.334,0.408)}{0.037}$ $\underset{(-0.320,0.372)}{0.026}$
BpAltStarsA [$\lambda=5$] $\mathbf{\underset{(-2.963,-2.390)}{-2.676}}$ $\mathbf{\underset{(-3.619,-3.010)}{-3.315}}$ $\mathbf{\underset{(-3.612,-3.006)}{-3.309}}$
BpAltStarsB [$\lambda=5$] $\mathbf{\underset{(0.408,0.524)}{0.466}}$ $\mathbf{\underset{(0.286,0.412)}{0.349}}$ $\mathbf{\underset{(0.279,0.406)}{0.343}}$
BpActivityA female $\underset{(-0.074,0.176)}{0.051}$ $\underset{(-0.060,0.100)}{0.020}$ $\mathbf{\underset{(-0.807,-0.300)}{-0.553}}$
BpContinuousActivityA age $\mathbf{\underset{(0.002,0.008)}{0.005}}$ $\mathbf{\underset{(0.001,0.007)}{0.004}}$ $\mathbf{\underset{(0.002,0.008)}{0.005}}$
BpActivityB industry.Personal.Goods $\mathbf{\underset{(0.103,0.314)}{0.208}}$ $\mathbf{\underset{(0.121,0.353)}{0.237}}$ $\mathbf{\underset{(0.109,0.330)}{0.219}}$
BpActivityB sector.Oil.and.Gas $\underset{(-0.031,0.008)}{-0.012}$ $\mathbf{\underset{(-0.059,-0.006)}{-0.032}}$ $\mathbf{\underset{(-0.060,-0.006)}{-0.033}}$
BinaryPairInteraction gender.Female sector.Oil.and.Gas $\underset{(-0.237,0.034)}{-0.101}$ $\underset{(-0.198,0.020)}{-0.089}$ $\underset{(-0.166,0.020)}{-0.073}$
BinaryPairInteraction gender.Male industry.Personal.Goods $\mathbf{\underset{(-0.553,-0.221)}{-0.387}}$ $\mathbf{\underset{(-0.573,-0.219)}{-0.396}}$ $\mathbf{\underset{(-0.543,-0.201)}{-0.372}}$
BpNodematchB country [$\beta=0.1$] $\mathbf{\underset{(2.231,2.794)}{2.513}}$ $\mathbf{\underset{(2.220,2.804)}{2.512}}$
BpNodematchB industry [$\beta=0.1$] $\mathbf{\underset{(0.765,1.408)}{1.087}}$ $\mathbf{\underset{(0.756,1.417)}{1.087}}$
BpNodematchB sector [$\beta=0.1$] $\mathbf{\underset{(0.148,0.507)}{0.328}}$ $\mathbf{\underset{(0.147,0.509)}{0.328}}$
BpNodematchA gender [$\beta=0.1$] $\mathbf{\underset{(0.146,0.486)}{0.316}}$
BpNodematchA female [$\beta=0.1$] $\mathbf{\underset{(0.731,1.587)}{1.159}}$
BpNodematchA male [$\beta=0.1$] $\underset{(-0.030,0.268)}{0.119}$
Just as for the Norwegian (Section 7.4) and Australian (Section 7.5) director interlock networks, the BipartiteAltStarsA parameter is negative and statistically significant, confirming a tendency against centralisation in the director degree distribution (number of boards a director sits on). In this network however, the BipartiteAltStarsB parameter is positive and statistically significant, implying that there is centralisation in the board degree distribution (boards that have large numbers of directors).
Goodness-of-fit plots for Model 3 are shown in Fig. D6. The model fits the degree distributions well, but not the other statistics (giant component size, dyadwise shared partner distribution, and cycle length distribution; the network is too large to test the fit to the geodesic cycle length distribution in reasonable time). Specifically, the model generates significantly fewer four-cycles than the observed network has, which is perhaps not surprising, as the model contains no terms to model four-cycles. This therefore shows that the international director interlock network contains significantly more four-cycles than expected, given the other effects included in the model, which includes terms for density, centralisation on degree, and gender, industry, and homophily effects.
Model 1 is no different from the model in Stivala et al. [100]. The statistically significant effects in this model confirm that older directors tend to be on more boards, companies in the personal goods industry tend to have larger boards, and men are less likely to be on boards in the personal goods industry. The corresponding tests for women in the oil and gas industry are not statistically significant.
All of the homophily effects added in Model 2 are positive and statistically significant. Companies in the same country are more likely to share directors than those that are not, as are companies in the same industry, and directors of the same gender are more likely to be on the same boards. Model 3 adds terms for differential gender homophily, and this effect is positive and statistically significant for female but not male directors.
7.8 Summary
Table 11 summarises the models estimated for the empirical networks. No model containing the FourCycles parameter could be estimated for any of the networks, and neither could any model containing the BipartiteAltKCyclesA parameter (and so no models containing both BipartiteAltKCycles parameters could be estimated). For four of the seven networks, a model containing BipartiteAltKCyclesB could be estimated. In contrast, for all four networks where models using the new BipartiteFourCyclesNodePower parameters could be estimated, both BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters were included in the model. Although both methods (BipartiteAltKCycles and BipartiteFourCyclesNodePower) could be used in only four of the seven networks, they are not the same four: there is one case (Scientific collaborations, Section 7.6) where BipartiteAltKCycles could be used but not BipartiteFourCyclesNodePower, and one case where the reverse is true (St Louis crime, Section 7.2).
Table 11: Summary of models fit to empirical networks. “neg” and “pos” indicate negative and positive parameter estimates, respectively, with * indicating statistical significance, while“n.c.“ indicates that no converged model could be found.
Network BipartiteAltKCycles BipartiteFourCyclesNodePower
A B A B
Inouye-Pyke pollinators n.c. pos neg pos*
St Louis crime n.c. n.c. pos pos
Norwegian director interlock n.c. pos pos* neg
Robertson pollinators n.c. n.c. n.c. n.c.
Australian director interlock n.c. neg pos* pos
Scientific collaborations n.c. neg* n.c. n.c.
International director interlock n.c. n.c. n.c. n.c.
8 Conclusion
The existing parameters for modelling shared partners or four-cycles in ERGMs for two-mode networks frequently lead to convergence problems, especially when parameters for both modes are included in the model. A literature survey shows that, perhaps as a result, the majority of published ERGM models of two-mode networks do not include these parameters. In addition, the majority of published models do not include an assessment goodness-of-fit to four-cycles.
In this work, we defined new ERGM effects to explicitly model four-cycles, and in the case of two-mode networks, four-cycle counts for the two modes separately. Simulation experiments show that ERGMs using the new parameters are able to generate bipartite networks with smoothly varying numbers of four-cycles, where the existing parametrisations cannot.
The empirical value of the new parametrisation was demonstrated on seven empirical two-mode networks, ranging in size from hundreds of nodes to hundreds of thousands of nodes. Four of the seven networks could have models estimated with existing parameters, but in all four cases only the parameter for one of the two modes could be included. In all four cases where converged models could be estimated with the new parameters, the parameters for both modes could be included, and the fit to the four-cycle count was better than that for models using existing ERGM parameters, where these could be estimated. We failed to estimate ERGM models at all for two of the networks, and for one network a model could only be estimated using the existing parameters, and not with the new ones. The new parameters are therefore not a complete solution to the problem of estimating ERGM models of two-mode networks, but are complementary to existing parameters, and further work remains in improving the ability to fit four-cycles (and longer cycles) in bipartite networks with ERGM models.
These new parameters come with both conceptual and computational costs. The conceptual cost is having to use a more general dependence class in the dependence hierarchy, without an underlying theoretical justification. The social circuit dependence assumption has theoretical justifications, as does partial inclusion ($\mathrm{PI}_{1}$) for pendant-triangle configurations [77], but we are forced to use the more general class $\mathrm{D}_{2}$ simply by the dependency induced by the new configuration, without any theoretical basis. The computational cost is incurred by having to traverse the two-neighbourhood of a dyad in computing the change statistic, and even with precomputation of two-path counts [71], this can be prohibitively slow (or require impractical amounts of memory) on dense and/or high four-cycle count or very large networks, or those with very high degree nodes, such as the Robertson pollinator network and international director interlock network.
Another shortcoming is the counterintuitive and possibly confusing interpretation of the new parameters for bipartite networks (BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB; see Section 6.1). The new parameter for one-mode networks (FourCyclesNodePower) is also applicable to two-mode networks but does not treat the modes separately, and is relatively straightforward to interpret simply as a four-cycle closure parameter, so one solution is to use this, but this could omit important differences in the modes in a two mode network. Even the relatively straightforward gwdegree parameter used in statnet causes significant confusion due to the counterintuitive meaning of its sign [106, 107, 75, 108]. Having to use such relatively complex statistics, such as alternating, geometrically weighted, or the “$\alpha$-outside” or “$\alpha$-inside” [76] weightings often solves the problem of near-degeneracy, but comes at the cost of making the statistics difficult to interpret [35, p. 400]. An alternative solution to the conceptual (and possibly computational) problem is to use instead the LOLOG model [60, 59] or tapered ERGM [31, 35, 98] where a simple four-cycle parameter is unlikely to cause near-degeneracy problems, much as they enable a simple triangle parameter in one-mode networks where alternating or geometrically weighted parameters are required in standard ERGMs [59, 35, 97]. LOLOG, however, does not (currently) handle bipartite networks [59, 109], while tapered ERGM has the advantage of being implemented in the statnet framework and can use any terms in the statnet ergm package.
One further shortcoming of the new parameters is that the weighting parameter $\alpha$ is fixed, and not estimated as part of the model. A potential avenue of future work is to explore the possibility of estimating this parameter in the context of a curved ERGM in the statnet ergm package.
One final potential avenue for future work is in fitting cycles of length larger than four in bipartite networks. It appears that sometimes fitting four-cycles better also fits longer cycles better, but sometimes it does not (Fig. 7). Of particular importance are six-cycles, which have been suggested as the basis (rather than four-cycles) for measuring closure in two-mode networks [10]. The dependence class required for the new statistics described in this work also admits six-cycles as configurations, however we did not attempt to fit models with six-cycles as a parameter. It seems likely that, like the simple four-cycles parameter, attempting to do so would lead to problems with near-degeneracy, necessitating the creation of another weighted configuration analogous to those defined here for four-cycles (which would then be in another, even more general, dependence class).
Funding
This work was supported by the Swiss National Science Foundation [grant number 200778].
Acknowledgements
This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.
Appendix A Models contained in literature survey
The list of 117 models in 59 publications is shown in Table LABEL:tab:literature. This table shows, for each network for which one or more models are presented in a paper, the size of both node sets in the network ($N_{A}$ is the number of nodes in node set $A$, and $N_{B}$ the number of nodes in node set $B$), the method used to estimate the model(s) in the paper, and the number of models for the network. Note that, in the case of multiple models for the same data, only a single model is counted; this is commonly the case for a model shown as developed incrementally, from a simple baseline model, and including more effects in subsequent models. In such cases, we consider only the final model. Multiple models are counted for cases such as models estimated for the same network at different time points, or for the same node set but with different edge types.
Table A1: Models included in literature survey of bipartite ERGM applications
Network size Estim. Num.
Citation Network description $N_{A}$ $N_{B}$ method models
[110] Theatregoers and theatre performances 290 24 MPLE 1
[54] Theatregoers and theatre performances 290 24 MPLE 1
[53] Municipalities and public utilities 116 120 Bergm 1
[55] Shareholder activists and firms 162 220 statnet 1
[111] Organizations and projects 198 95 BPNet 4
[112] Customers and products (cars) 5000 250 statnet 6
[113] Affiliation network (club membership) 257 15 BPNet 1
[114] Actors and issues in water policy ? ? statnet 1
[115] Software contributors and software bugs 72 737 BPNet 1
[116] Lemur species and plant genera 55 590 statnet 1
[117] Buyers and sellers in an online drug cryptomarket 706 57 statnet 1
[118] Buyers and sellers in an online drug cryptomarket 706 57 statnet 1
[119] Co-attendance of Soviet politicians at events 67 1816 MPLE 8
[120] Climate change adaptation actors and issues 659 19 statnet 1
[121] Patents and inventors (temporal) 78 412 126 388 statnet 1
[122] Fruit-frugivore interactions (vertebrates and plants), Continuum 133 315 statnet 1
[122] Fruit-frugivore interactions (vertebrates and plants), Fragment 54 58 statnet 1
[123] Provincial and national organizations 29 52 statnet 1
[124] Authors and papers 2 200 76 BPNet 1
[125] University departments and specializations 101 77 MPNet 1
[126] Policy actors and decision-making forums 109 84 statnet 1
[72] Director interlock, largest 248 corporations by revenue in Australia 1 251 248 BPNet 1
[127] Organizations and research projects 1 316 237 BPNet 2
[128] Interest groups and lobbying coalitions 171 74 statnet 1
[129] Characters and their groups in an online game 465 396 BPNet 7
[130] Organizations and forums in watershed policy ? ? statnet 1
[131] Organizations and policy institutions 527 146 BPNet 1
[132] Wikipedia articles and editors 14 292 249 statnet 1
[133] Southern Women 18 14 statnet 1
[133] World city network (global firms and cities) 100 315 statnet 1
[133] Inventors and patents 10 251 8 206 statnet 1
[134] Travel destination countries and attitudes towards them 47 22 statnet 1
[135] Disaster response organizations and resource types 27 10 BPNet 3
[135] Affected neighbourhoods and resource contacts 19 10 BPNet 4
[136] EU parliamentary chambers and proposals 39 650 statnet 1
[137] Collaborations for hazard mitigation 95 198 statnet 1
[138] Online doctor-patient consultations ? ? statnet 1
[139] Actors and institutions in water management 167 220 BPNet 2
[56] Actors and forums in adapting to sea-level rise 82 9 statnet 1
[56] Actors and forums in adapting to sea-level rise 647 103 statnet 1
[140] Voluntary collaborative project teams 170 124 BPNet 1
[141] Aged care residents and group activities 35 563 statnet 1
[142] Organizations and policy forums 52 16 BPNet 2
[143] Stakeholders and committees/working groups 152 9 BPNet 2
[143] Stakeholders and committees/working groups 180 9 BPNet 5
[144] Stakeholders and forums in urban development 60 14 MPNet 1
[145] Stakeholders and forums in myrtle rust response 259 12 MPNet 1
[146] Actors and instruments in water policy 31 15 statnet 1
[147] "Swinging" couples and venues 57 39 BPNet 1
[148] Organizations and projects in EU conservation (UK) 120 46 BPNet 1
[148] Organizations and projects in EU conservation (NL) 54 40 BPNet 1
[148] Organizations and projects in EU conservation (PT) 156 63 BPNet 1
[148] Organizations and projects in EU conservation (GR) 124 57 BPNet 1
[148] Organizations and projects in EU conservation (RO) 119 48 BPNet 1
[148] Organizations and projects in EU conservation (LV) 122 28 BPNet 1
[149] Buyers and sellers in an online drug cryptomarket 3 542 463 statnet 1
[150] Citations from web top-level domains to health agency websites ? 148 statnet 1
[151] Stakeholders and resilience planning documents ? 39 statnet 1
[152] Drivers and regime shifts in social-ecological systems 57 25 statnet 1
[153] Actors and forums in collaborative governance regimes 400 57 statnet 3
[154] Customers and car models 5 000 281 statnet 1
[155] Participants in design crowdsourcing contests 3 462 96 statnet 1
[156] Deforestation emissions reduction projects and countries 480 57 MPNet 1
[157] Crime control bills and their sponsors 1 304 221 statnet 1
[158] Employees and ideas in intraorganizational crowdsourcing (stage 1) 213 236 BPNet 3
[158] Employees and ideas in intraorganizational crowdsourcing (stage 2) 685 578 BPNet 3
[158] Employees and ideas in intraorganizational crowdsourcing (full) 768 640 BPNet 3
[18] Southern Women 18 14 BPNet 1
[18] Director interlock 366 50 BPNet 1
[18] Director interlock 255 198 BPNet 1
[19] Student activists youth leaders by event 14 49 BPNet 1
[19] Student activists organization by event 23 49 BPNet 1
[159] Participants and teams in online crowdsourcing 2 100 946 statnet 1
[160] EU Parliamentarians and information sources 77 18 statnet 1
[161] Individuals and teams in an online game 333 426 BPNet 1
Appendix B A simpler, but unsuccessful, new statistic
Wang et al. [18] define the alternating $k$-two-path statistics $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$, implemented as XACA and XACB in MPNet [44, 45], and BipartiteAltKCyclesA and BipartiteAltKCyclesB in EstimNetDirected (https://github.com/stivalaa/EstimNetDirected):
$z_{\text{BipartiteAltKCyclesA}(\lambda)}=\lambda\sum_{i\in B}\sum_{\left\{l\in B\,:\,l<i\right\}}\left[1-\left(1-\frac{1}{\lambda}\right)^{L_{2}(i,l)}\right]$ (22)
where $\lambda>1$ is the decay parameter, and $L_{2}(i,l)$ is the number of two-paths connecting $i$ to $l$, as defined by equation (2) in the main text. The corresponding change statistic is [18]:
$\delta_{\text{BipartiteAltKCyclesA}(\lambda)}(i,j)$ $=\sum_{l\in B}\left[x_{il}\left(1-\frac{1}{\lambda}\right)^{L_{2}(j,l)}\right]$ (23)
$=\sum_{l\in N(i)}\left(1-\frac{1}{\lambda}\right)^{L_{2}(j,l)}.$ (24)
The statistic and change statistic for BipartiteAltKCyclesB are defined similarly.
As discussed in the main text (Section 4), one problem with this statistic, particularly if used to model bipartite closure, is that it counts two-paths (as its highest-weighted term) and not just four-cycles, and hence open paths, or cycles of length greater than four, contribute to the statistic, even if there are no four-cycles at all. A simple solution to this problem, therefore, is to remove the first term and reverse the signs, so that it no longer counts open two-paths, but the first, positive, term actually counts four-cycles. We therefore define a new statistic BipartiteAltK4CyclesA as
$z_{\text{BipartiteAltK4CyclesA}(\lambda)}=-\left(z_{\text{BipartiteAltKCyclesA}(\lambda)}-z_{\text{TwoPathsA}}\right),$ (25)
and its change statistic
$\delta_{\text{BipartiteAltK4CyclesA}(\lambda)}(i,j)=-\left(\delta_{\text{BipartiteAltKCyclesA}(\lambda)}-\operatorname{deg}(i)\right)$ (26)
where
$z_{\text{TwoPathsA}}=\sum_{i\in B}\sum_{\left\{l\in B\,:\,l<i\right\}}L_{2}(i,l)$ (27)
is the number of two-paths connecting nodes in node set $B$ (and which therefore go through a node in node set $A$), and $\operatorname{deg}(i)$, the degree of node $i$, is the change statistic for the number of two-paths through node $i$. BipartiteAltK4CyclesB and its change statistic are defined similarly. Unlike the new FourCyclesNodePower statistics defined in the main text (Section 5), these new statistics are in the same dependence class ($\mathrm{I}_{1}$, “social circuit”) as the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics from which they are derived.
Table B1 is a copy of Table 3 in the main text, but with the new BipartieAltK4CyclesA and BipartiteAltK4CyclesB statistics (labelled BpAK4CA and BpAK4CB respectively) included (and the $\mathrm{N_{A}}$, $\mathrm{N_{B}}$, L, BpNP4CA, and BpNP4CB columns removed to make space). This table shows the value of the new statistics on some small example networks, demonstrating, that, by design, they are zero for networks in which there are no four-cycles.
Table B1: Statistics of some example bipartite networks. C4 is the number of four-cycles,and BpAK4CA and BpAK4CB are the new statistics BipartiteAltK4CyclesA and BipartiteAltK4CyclesB, respectively. Nodes in node set A are represented as red circles, and nodes in node set B as blue squares.
Name Visualization C4 XACA XACB BpAK4CA BpAK4CB
Two-path [Uncaptioned image] 0 1 0 0 0
Four-cycle [Uncaptioned image] 1 1.5 1.5 0.5 0.5
Four-cycles-3 [Uncaptioned image] 3 4.5 1.75 1.5 1.25
Ten-cycle [Uncaptioned image] 0 5 5 0 0
Nine-star [Uncaptioned image] 0 36 0 0 0
Four-fan-3 [Uncaptioned image] 3 16.5 4.5 1.5 1.5
Unfortunately, however, simulation experiments indicate that the new BipartiteAltK4CyclesA and BipartiteAltK4CyclesB parameters are actually more problematic with respect to near-degeneracy than the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters. Fig. B1 shows the results of simulation experiments similar to those described by Wang et al. [18, p. 19]. Bipartite networks were simulated with 30 nodes in node set $A$ and 20 nodes in node set $B$ with the Edge parameter set to $-3.0$. In three different sets of simulations, for each of the BipartiteAltKCyclesB, BipartiteAltK4CyclesB, and BipartiteFourCyclesNodePowerB parameters, the parameter in question is varied from $-1.00$ to $10.0$ in increments of $0.01$ for each of two values of $\lambda$ ($\lambda=2$ and $\lambda=5$) for BipartiteAltKCyclesB and BipartiteAltK4CyclesB, and for each of two values of $\alpha$ ($\alpha=1/2$ and $\alpha=1/5$) for BipartiteFourCyclesNodePowerB. The networks were simulated using the SimulateERGM program from the EstimNetDirected software package, using the basic ERGM sampler, with a burn-in of $10^{5}$ iterations and an interval of $10^{4}$ iterations between each of 100 samples, to ensure that samples are drawn from the equilibrium ERGM distribution, and are not too autocorrelated.
Figure B1: Effect of varying the BipartiteAltKCyclesB parameter (top), BipartiteAltK4CyclesB parameter (middle), and BipartiteFourCyclesNodePowerB parameter (bottom) on the statistic corresponding to the parameter itself (left) and the Edge statistic (right). Each graph plots the mean value of the statistic (over 100 simulations) for two different values of the relevant $\lambda$ or $\alpha$ parameter.
The top right panel in Fig. B1 shows the results of a simulation similar to that shown in Fig, 9 of Wang et al. [18]: the number of edges increases smoothly, giving good coverage of the graph space (with respect to density, at least). The top left panel shows the value of the statistic corresponding to the parameter BipartiteAltKCyclesB itself. The behaviour of this curve is starting to look as if it could be prone to near-degeneracy, with a fairly steep increase at a critical value.
The graphs in the middle row of Fig. B1 show the results for the new BipartiteAltKCyclesB parameter. In this case, both the Edge statistic (right) and BipartiteAltK4CyclesB statistic itself (left) show a phase transition, where a critical value of the parameter separates an empty graph regime from a complete graph regime. This model is therefore near-degenerate, suggesting that this new parameter may in fact be less useful than the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters.
For completeness, the bottom row of Fig. B1 shows the results for the new BipartiteFourCyclesNodePowerB parameter defined in the main text (Section 5). When $\alpha=0.5$ an abrupt change from a near-empty to a full graph occurs, however, as shown by the result for $\alpha=0.2$, this can be removed by decreasing the value of $\alpha$.
Perhaps unsurprisingly, given the near-degeneracy demonstrated in Fig. B1, we were able to find a converged model with the new parameter for only one of the seven empirical networks described in the main text (Section 7). This is the smallest of the seven, the Inouye-Pyke pollinator web. Table B2 shows this model as Model 3 (Models 1 and 2 reproduce the models shown in Table 5 in the main text). Models with BipartiteAltK4CyclesA did not converge, and the value of $\lambda$ for BipartiteAltK4CyclesB had to be reduced from the default ($\lambda=2$) while the value of $\lambda$ for BipartiteAltStarsB had to be increased in order for this model to converge.
Table B2: ERGM parameter estimates for the Inouye-Pyke pollinator web. including Model 3 with the BipartiteAltK4Cycles parameter, with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”.
Effect Model 1 Model 2 Model 3
Edge $\mathbf{\underset{(-6.366,-5.598)}{-5.982}}$ $\mathbf{\underset{(-6.470,-5.989)}{-6.229}}$ $\mathbf{\underset{(-6.012,-5.416)}{-5.714}}$
BpAltStarsA [$\lambda=2$] $\underset{(-0.552,2.397)}{0.923}$ $\underset{(-1.674,0.878)}{-0.398}$
BpAltStarsA [$\lambda=5$] $\mathbf{\underset{(0.029,0.965)}{0.497}}$
BpAltStarsB [$\lambda=5$] $\underset{(-0.493,1.597)}{0.552}$ $\mathbf{\underset{(0.319,1.706)}{1.013}}$ $\underset{(-0.024,1.014)}{0.495}$
BpAltKCyclesB [$\lambda=2$] $\underset{(-0.322,0.332)}{0.005}$
BpFourCyclesNodePowerA [$\alpha=1/5$] $\underset{(-4.443,1.662)}{-1.391}$
BpFourCyclesNodePowerB [$\alpha=1/5$] $\mathbf{\underset{(1.670,11.401)}{6.535}}$
BpAltK4CyclesB [$\lambda=1.1$] $\mathbf{\underset{(0.004,0.102)}{0.053}}$
The interpretation of Model 3 (Table B2) is similar to that of Model 1 and Model 2 discussed in Section 7.1. The positive and significant BipartiteAltStarsA parameter indicates centralisation on pollinator nodes, that is, the presence of pollinators that pollinate many plants. The positive and significant BipartiteAltK4CyclesB parameter indicates an over-representation of pollinators (node set $A$) participating in multiple four-cycles.
Goodness-of-fit plots for Model 3 are shown in Fig B2. Compared to the goodness-of-fit for Model 2 (Fig D1), the fit to the degree distribution for mode B nodes is worse, as is the fit to the geodesic distance distribution, however the fit to the cycle length distribution is slightly better, in particular for cycles of length 12.
Figure B2: Goodness-of-fit plots for Inouye-Pyke pollinator web ERGM Model 3 (Table B2). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
In summary, we defined new BipartiteAltK4CyclesA and BipartiteAltK4CyclesB parameters as modified forms of the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters defined in Wang et al. [18], in order to count four-cycles but not open two-paths. However, near-degeneracy was exhibited in the simulation experiments (similar results occur with the larger simulated networks described in Section 6; data not shown), and a converged model using them could only be found for one of the seven networks tested. We conclude, therefore, that although these new parameters could potentially be useful in some cases, they are more prone to near-degeneracy and not as useful as either the existing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters, or the new BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters described in the main text.
Appendix C Statnet implementation example
The new statistics BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB were implemented in statnet as b1np4c and b2np4c, using the facility to define custom ergm model terms [80, 81]. They are available at https://github.com/stivalaa/ergm.terms.contrib.
Here we demonstrate this implementation by using the the statnet ergm package [48] with the new user terms to estimate a model for the Davis “Southern Women” network [162], obtained via the latentnet R package [163, 164]. This well-known affiliation network represents the participation of 18 women (first mode) in 14 social events (second mode).
The b1np4c and b2np4c model terms take the $\alpha$ ($0<\alpha\leq 1$) value as a parameter, with a default value of $\alpha=0.5$ if omitted. For example, to estimate a model with the b2np4c (BipartiteFourCyclesNodePowerB) term with $\alpha=1/5$ (Model 4 in Table C1):
davis_model4 <- ergm(davis ~ edges + gwb1degree(1, TRUE) +
                             gwb2degree(1, TRUE) + b2np4c(1/5),
  control = control.ergm(main.method = "Stochastic-Approximation"))
Table C1: ERGM parameter estimates for the Southern Women network, estimated with statnet. The table was generated directly from the statnet models with the texreg R package [165].
Model 1 Model 2 Model 3 Model 4
edges $-2.07^{***}$ $-0.20$ $0.47$ $-5.90^{***}$
$(0.34)$ $(0.24)$ $(0.31)$ $(0.58)$
b1star2 $0.07$
$(0.07)$
b2star2 $0.18^{***}$
$(0.04)$
gwb1deg.fixed.1 $-0.83$ $-7.04^{***}$ $10.60^{***}$
$(1.03)$ $(1.51)$ $(1.80)$
gwb2deg.fixed.1 $-2.26^{**}$ $7.76^{**}$ $-6.95^{***}$
$(0.85)$ $(2.48)$ $(1.57)$
gwb1dsp.fixed.0.5 $0.45^{***}$
$(0.12)$
gwb2dsp.fixed.0.5 $-1.33^{***}$
$(0.30)$
b2np4c.fixed.0.2 $17.30^{***}$
$(1.86)$
AIC $319.34$ $328.85$ $308.77$ $285.41$
BIC $329.93$ $339.44$ $326.41$ $299.53$
Log Likelihood $-156.67$ $-161.43$ $-149.38$ $-138.71$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
Table C1 shows four models for the Southern Women network, estimated with the stochastic approximation algorithm [24]. Model 1, with only the edges and two-star terms for each mode, is the same as [18, Model (8.3)]. Model 2, using the geometrically weighted degree terms rather than two-stars, is similar to [18, Model (8.5)], but using the statnet gwb1degree and gwb2degree terms rather than the BPNet alternating $k$-star terms $\mathrm{K\text{-}S_{P}}$ and $\mathrm{K\text{-}S_{A}}$. Note the reversal of interpretation of signs between $\mathrm{K\text{-}S_{P}}$/$\mathrm{K\text{-}S_{A}}$ and gwb1degree/gwb2degree [106, 107]. Model 3, adding the geometrically weighted dyadwise shared partner terms, is similar to [18, Model (8.6)]. Model 4 uses the new b2np4c term rather than the geometrically weighted dyadwise shared partner terms gwb1dsp and gwb2dsp (models with b1np4c did not converge).
Cycle length distribution goodness-of-fit plots for the four models are shown in Fig. C1, and statnet goodness-of-fit plots in Fig. C2. Note that all four models fit acceptably well on all the statistics included in the goodness-of-fit tests (degree distributions for each mode, dyadwise shared partners, and geodesic distance distribution), as well as the cycle length distributions. In particular, the fit to four-cycle counts is good for all models, including Model 1 and Model 2, which do not contain any terms to model four-cycles. It therefore appears that Model 1 is the most parsimonious explanation of this data, just as discussed in Wang et al. [18, pp. 22]. This model has a positive and statistically significant event two-star parameter (b2star2), indicating “greater discrepancies in the popularity of events than expected in a random network” [18, pp. 22], taking into account the density (edges) and actor two-star (b1star2) effects.
Figure C1: Cycle length distribution goodness-of-fit plots for the Southern Women network ERGM (Table C1) Model 1 (top left), Model 2 (top right), Model 3 (bottom left), and Model 4 (bottom right). Observed network statistics are plotted as red points (joined by red lines as a visual aid) with the statistics of 100 simulated networks plotted as black box plots.
Figure C2: Statnet goodness-of-fit plots for the Southern Women ERGM (Table C1) Model 1 (top left), Model 2 (top right), Model 3 (bottom left), and Model 4 (bottom right).
That ERGM models containing only terms to model density and degree distributions (and specifically not any terms to model dyadwise shared partner distributions or four-cycles) also fit the four-cycle count well indicates that the observed number of four-cycles could have occurred simply by chance [18, pp. 22]. Note that the same applies also to six-cycles. This is consistent with the results for this network described by Opsahl [10], where the observed value of the two-mode global clustering coefficient defined in that paper is not extreme in the distribution of that coefficient in random networks.
Appendix D Goodness-of-fit plots
Goodness-of-fit plots for the ERGM models for the Inouye-Pyke pollinator web (Table 5), the St Louis crime network (Table 6), the Norwegian director interlock network (Table 7), the Australian director interlock network (Table 8), the scientific collaboration network (Table 9), and the international director interlock network (Table 10) are shown in Fig. D1, Fig. D2, Fig. D3, Fig. D4, Fig. D5, and Fig. D6, respectively.
Figure D1: Goodness-of-fit plots for Inouye-Pyke pollinator web ERGM (Table 5) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D2: Goodness-of-fit plots for the St Louis crime network ERGM (Table 6) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D3: Goodness-of-fit plots for the Norwegian director interlock network ERGM (Table 7) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D4: Goodness-of-fit plots for the Australian director interlock network ERGM (Table 8) Model 2 (top) and Model 3 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D5: Goodness-of-fit plots for the scientific collaboration network ERGM (Table 9) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D6: Goodness-of-fit plots for the international director interlock network ERGM (Table 10) Model 3. The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Appendix E Supplementary Tables
Table E1: Four-cycle descriptive statistics.
Network $\mathrm{N_{A}}$ $\mathrm{N_{B}}$ 4-cycles Unique nodes
in 4-cycles
A B
Inouye-Pyke pollinators 91 42 830 47 33
St Louis crime 870 557 144 93 121
Norwegian director interlock 1 495 367 108 89 77
Robertson pollinators 1 428 456 565 969 862 456
Australian director interlock 9 971 2 087 1 743 742 663
Scientific collaborations 16 726 22 015 70 549 7 604 14 137
International director interlock 321 869 34 769 79 885 12 567 8 136
Table E2: ERGM parameter estimates from MPNet for the Inouye-Pyke pollinator web. * indicates statistical significance at the nominal 95% level.
Effects Lambda Parameter Stderr t-ratio
XEdge 2.$0$5.5726 0.197 0.024 *
XASA 2.0000 0.7998 0.127 0.022 *
XASB 2.0000 0.7301 0.174 0.020 *
XACB 2.0000 0.0730 0.002 -0.012 *